### v2

Abstract
The full family of discrete logistic maps has been widely studied both as a canonical example of the period-doubling route to chaos, and as a model of natural processes. In this paper we present a study of the stochastic process described by iterations of the logistic map with a random parameter. In addition to bringing together previously known results, we present a proof that the unique invariant measure of the process in the chaotic regime is asymptotically stable.
keywords:
stochastic quadratic map; discrete chaos; Markov process; asymptotic stability
1 Introduction
The discrete logistic map:
$f_{\lambda}(x)=\lambda x(1-x)$,
is well known as a model of self-limiting population growth, whose dynamic properties are described in the early review paper by Robert May [1]. In this highly influential paper (over 6,600 citations as of November, 2023), May states that, while the fine structure of the bifurcation diagrams in unimodal maps are fascinating, what is more important in practice is the statistical properties of these maps. In his opinion, “What is called for is an effectively stochastic description of the dynamics”, and that these descriptions of evolving continuous densities will still capture some salient features of the maps themselves. The bifurcation diagram of the logistic family shows how the long term dynamics of the map change as the parameter, $\lambda$, is varied. What May is referring to in this challenge is a description of the distribution of $x$-values, the “continuously evolving densities”, as the parameter $\lambda$ varies stochastically. In practice, we might want to model a scenario in which the intrinsic growth rate of a population fluctuates over a specific range by letting the parameter $\lambda$ be a random variable with support in a small interval.
In the literature on random dynamical systems, a stochastically-varying parameter is known as “parametric noise” to distinguish it from “additive noise”. This distinction is important from a modeling perspective: parametric noise models stochastic features of the environment, be it reproduction rates that vary due to temperature or resources, or economic growth rates, which might vary in response to world events, politics, or other economic pressures. Additive noise, on the other hand, models uncertainty in the model predictions, so that the model becomes: $X_{n+1}=f(X_{n})+\epsilon_{n}$, where $\epsilon_{n}$ is the additive noise which is typically independent of the previous observation, $X_{n}$, and of the law of evolution, $f(\cdot)$. In this paper we consider parametric noise, modeling uncertainty in the model parameters. Mathematically, parametric noise gives rise to some interesting questions, such as the uniqueness and stability of invariant measures. We note that, in some cases, additive noise which is state-dependent (i.e. $\epsilon_{n}$ depends on $X_{n}$) can be reinterpreted as parametric noise, so the distinction between the two is not always clear. See, for example, [2] which studies quasi-stationary distributions of the logistic map with state-dependent additive noise. For a theoretical discussion of additive noise in the context of random dynamical systems, we refer the reader to [3]. The logistic map with additive noise has also been studied in physical and computational contexts. For example, in [4] this type of random logistic map been studied as a model of physical processes with random “kicks”. Other examples (a far from complete list!) include [5], which discusses how bifurcations of the logistic map are affected by additive noise, [6], which describes how synchronization time can be shortened with additive noise, and [7] which studies the random logistic map with additive noise in the context of algorithmic probability.
Since May’s remarks appeared, many models involving parametric noise have appeared in the ecology, biology and economics literature. In [8], the authors address the Neubert conjecture by looking at a stochastically-perturbed Beverton-Holt model of population growth. A continuous version of the stochastic logistic model is studied in [9], where they show that these systems can be used to describe experimental results on microbial growth. Three different economic models described by randomly perturbed deterministic maps are studied in [10], [11] and [12]. These examples are by no means exhaustive, but they do give a sense of the ubiquity and impact of these models. In most cases, the goal is to first show that the map has a unique invariant measure, since the uniqueness of this measure is essential if one is to say something about the long-term behavior of the system.
Along these same lines, Kolmogorov suggested starting with a deterministic dynamical system, and adding a stochastic perturbation in such a way that the resulting process has a unique invariant measure. This stochastic version of the original deterministic system is sometimes called the Kolmogorov measure of the dynamical system [13].
In this paper we present a study of the stochastic process described by iterations of the logistic map with a randomly-varying parameter, which we will refer to as “the stochastic logistic map”. Our contribution to the literature on the stochastic logistic map is that, in addition to bringing together previously known results, we present a proof that the unique invariant measure of the process is asymptotically stable.
2 Background and Definitions
In this section we present some technical background, notation, and definitions for the reader who is new to the concepts of stochastic dynamical systems, transition probabilities and invariant measures.
Given a state space $S$, one can consider either a deterministic or a stochastic (or random) discrete dynamical system on $S$. A deterministic dynamical system is given by a function $f:S\rightarrow S$, where, given an initial state $x_{0}\in S$, one can compute future states recursively:
$x_{n+1}=f(x_{n}).$
Because, given $x_{n}$, the value of $x_{n+1}$ can be computed exactly, this is a deterministic process, and the set of iterates, $x_{0},x_{1},x_{2},\dots=\{x_{n}\}_{n=0}^{\infty}$, is called the trajectory of the initial value $x_{0}$ under the dynamical system given by the function $f$. A natural question is: given an initial value, $x_{0}$, what is the long-term behavior of the sequence $x_{0},x_{1},x_{2},\dots$? If the sequence converges to a finite value, it must be a fixed point, i.e. a value $x^{*}$ that satisfies $x^{*}=f(x^{*})$. So one of the first questions a dynamicist asks when encountering a dynamical system is: where are the fixed points? Other questions of interest include determining the stability of fixed points - whether nearby trajectories converge toward, or away from, fixed points - and looking at the limiting behavior of non-fixed points.
In a stochastic dynamical system, we consider a sequence of random variables: $\{X_{n}\}_{n=0}^{\infty}$. Given a value for the random variable $X_{n}$ in the state space, the value of $X_{n+1}$ can not be determined exactly because the law of evolution is subject to some noise. We are then interested in the sequence of distributions of these random variables. In particular, we are interested in the existence of an invariant distribution: a distribution of $X_{n}$ such that $X_{n+1}$ has the same distribution. If such a distribution exists, we are also interested in its stability; that is, whether other distributions converge to this invariant distribution in some manner. We define the random logistic map and the notions of invariant measures and stability more precisely below.
In this paper, we consider the state space $S=[0,1]$ and the particular choice of parameter space $\Lambda=(3.87,4)\subset[0,4]$. With these choices, we are ensured that any point in $S$ is mapped to a point in $S$, and that all of the logistic maps are in the “chaotic regime”. This will be important in the proofs in Section 3.
Let $\mathcal{B}(S)$ denote the Borel sigma-algebra on $S$. We consider the logistic Markov process
$X_{n+1}=\lambda_{n+1}X_{n}(1-X_{n}),$ (1)
where $\{\lambda_{i}\}_{i\geq 1}$ are independent, $\Lambda$-valued random variables, independent of $X_{n}$ with identical distribution $Q$. The initial state $X_{0}$ is an $S$-valued random variable, independent of the $\lambda$’s. Let $A\subset S$. Then $\{X_{n}\}$ is an $S$-valued Markov process with a continuous state space, $S$, and transition probabilities
$\hbox{Prob}\{X_{1}\in A|X_{0}=x\}=p(x,A)=Q\left(\frac{A}{x(1-x)}\right).$
In words: the probability that $X_{1}$ is in a given set $A$, given $X_{0}=x$, is the probability of the set of $\lambda$-values that map $x$ to a point in $A$. The measure $p$ is called a (one-step) transition probability on the state space, $S$, and we can write $p(x,A)=\int_{S}{\bf 1}_{A}\,p(x,dy)$.
Similarly, we write the $n$th order transition probability as
$\hbox{Prob}\{X_{n}\in A|X_{0}=x\}=p^{n}(x,A).$
Suppose we have a distribution, $\mu$, on the state space, $S$. When we apply the process (1) to subsets of $S$, we induce a new distribution, $P\mu$ on the state space as follows. If $A\subset S$, then we define
$P\mu(A)=\int_{S}p(x,A)\mu(dx).$ (2)
So the operator $P$ maps probability measures on $S$ to probability measures on $S$. We can now define an invariant measure under the operator, $P$, which corresponds to a fixed point in the deterministic case.
Definition 1.
If $P\mu(A)=\mu(A)$ for all $A\in{\cal B}(S)$ then $\mu$ is an invariant measure under the operator $P$.
We can iterate the operator $P$ a positive number $m$ times, which we denote by $P^{m}$: if $X_{0}$ has distribution $\mu$ then $X_{m}$ has distribution $P^{m}\mu$. This leads us the analog to stability in the stochastic context.
Definition 2.
A sequence of probability measures $P^{n}(n\geq 1)$ on $S$ converges weakly to a probability measure $P$ on $S$ if, for any continuous, bounded function $h:S\mapsto S$,
[…]
In the deterministic case, we know that, if the iterates of $x_{0}$ under a continuous function, $f$, converge to a point, $x$, then $x$ must be a fixed point of $f$. This result follows from the continuity of $f$. We can get a similar result in the stochastic setting, but we first need to define a notion of continuity. Suppose we have a Markov process on $S$ with transition probabilities given by $p(x,dy)$.
Definition 3.
The transition probability $p(x,dy)$ is weakly continuous or has the Feller property if, for every continuous and bounded function $h:S\mapsto S$, the function
$x\mapsto\int_{S}h(y)p(x,dy)$
is continuous.
Using weak continuity we get the following result ([14], Theorem 11.1).
Proposition 1.
Suppose $p(x,dy)$ is a transition probability on $S$ with the Feller property.
  1. 1.
    If there is an $x\in S$, a sequence of integers $1\leq n_{1}<n_{2}<\dots$ and a probability measure $\pi_{x}$ on $S$ such that
    $\frac{1}{n_{k}}\sum_{m=0}^{n_{k}-1}p^{(m)}(x,dy)\hbox{ converges weakly to }\pi_{x}\hbox{ as }k\rightarrow\infty$
    then $\pi_{x}$ is an invariant measure for $p$.
  2. 2.
    If the limit above exists for all $x\in S$ and is independent of $x$ then $\pi=\pi_{x}$ is the unique invariant probability for $p$.
We can state this in terms of the operator $P$ given in Equation (2). Suppose a process has the Feller property and we have a measure $\mu$ on $S$. If we have weak convergence in terms of the average to a measure, $\pi$:
$\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{m=0}^{n-1}P^{m}\mu(A){\rightarrow}\pi(A)$
then $\pi$ is an invariant measure under the operator $P$, i.e. $P\pi(A)=\pi(A)$ for $A\in{\cal B}(S)$. Furthermore, if Condition 2 holds in Proposition 1, then this invariant measure is unique.
What we would like is the asymptotic stability of the (unique) invariant measure, $\pi$. In other words, for any probability measure $\mu$ on $S$, when is it the case that
$\sup_{A\in{\cal B}(S)}|P^{n}\mu(A)-\pi(A)|\rightarrow 0\quad\hbox{ as }n\rightarrow\infty?$
To answer this question, we need the notions of irreducibility and aperiodicity.
The following definitions of irreducibility and recurrence are from [14], Def. 9.2
Definition 4.
Let $\mathcal{B}(S)$ denote the Borel subsets of $S$.
A Markov process $\{X_{n}\}_{n=0}^{\infty}$ is Harris irreducible with reference measure $\nu$ (or $\nu$-irreducible for short) if, for every $x\in S$ and every $B\in\mathcal{B}(S)$ with $\nu(B)>0$, there is $n\geq 1$ such that $p^{n}(x,B)>0$.
In words: a Markov process is Harris irreducible if every element in the state space eventually gets mapped to any set of positive probability.
In the theorems that follow, we require that the process be Harris irreducible with respect to any $sigma$-finite measure: the particular measure is not important. Therefore, to make the language more concise, we sometimes just say “Harris irreducible” to mean “Harris irreducible with respect to some $\sigma$-finite measure.”
The Krylov-Bogolubov theorem [3] tells us that a stochastic dynamical system obtained from a continuous function on a compact subset of ${\mathbb{R}}^{d}$ with independent parametric noise always has an invariant distribution:
Theorem 1 (Krylov-Bogolubov).
A Feller continuous, stationary Markov process on a compact subset of ${\mathbb{R}}^{d}$ has an invariant distribution.
The theorem applies to more general settings, but this more specific version is all we need here. Harris irreducibility can then be used to show that this invariant distribution is unique.
Theorem 2 ([14], Theorem 2.C9.3).
If a Markov process is Harris irreducible with respect to some $\sigma$-finite measure $\nu$ on $S$ and has an invariant probability $\mu^{*}$, then $\mu^{*}$ is the unique invariant probability and, for any probability measure $\mu$ on S:
$\lim_{n\rightarrow\infty}\sup_{A\in\mathcal{B}(S)}\left|\frac{1}{n}\sum_{m=1}^{n}P^{m}\mu(A)-\mu^{*}(A)\right|=0$
Thus, if a Markov process is Harris irreducible, we know that any measure converges to a unique invariant measure, $\mu^{*}$, under the operator, $P$ “in the sense of Césaro”, i.e. the average of the sequence of measures conerges to $\mu^{*}$. To get asymptotic stability of the invariant meausure without this average, we need one more condition: aperiodicity.
Definition 5.
A Markov process is said to be strongly aperiodic if there is a set $A_{0}\subset S$ with $\nu(A_{0})=1$ and an $\alpha>0$ such that for all $x\in A_{0}$, $A$ a measurable subset of $S$, one has
$p(x,A)\geq\alpha\nu(A)$
Theorem 3 ([14], Theorem C9.2).
If a Markov process is Harris irreducible with respect to some $\sigma$-finite measure $\nu$ on $S$ and strongly aperiodic with invariant measure $\mu^{*}$ then, for any $\mu\in\mathcal{B}(S)$,
$\lim_{n\rightarrow\infty}\delta(P^{n}\mu,\mu^{*})=0$
where $\delta$ is the total variation distance
$\delta(P,Q)=\sup_{B\in\mathcal{B}}|P(B)-Q(B)|.$
As an example of the ideas presented in this section, let us consider a version of the stochastic logistic map as examined in [15]. In this paper, Bhattacharya and Rao consider the stochastic logistic map
$X_{n+1}=\lambda_{n+1}X_{n}(1-X_{n})$
where $\lambda$ is distributed according to a Bernoulli distribution, and takes two possible values, $\alpha$ and $\beta$ (where $\alpha<\beta$). Thus, the process switches randomly between two quadratic maps. Bhattacharya and Rao demonstrate that if $0\leq\alpha<\beta\leq 1$, then $\delta_{0}$, the Dirac measure supported on 0, is the unique invariant probability on [0,1]. If $1<\alpha<\beta\leq 2$ then there is a unique invariant probability on (0,1), and this probability is nonatomic. If $2<\alpha<\beta<1+\sqrt{5}$ and $\alpha\in[8/\beta(4-\beta),\beta)$, then there exists a unique invariant probability on (0,1) which is nonatomic and has its support contained in $[1/2,\lambda/4]$.
3 Proof of the asymptotic stability of the invariant measure in the chaotic regime.
In section 2, we outlined results that give conditions for the existence of an invariant distribution of the stochastic logistic map, and for the stability of this invariant distribution. In this section we use these results to prove that, when parameters are supported in the chaotic regime, the random logistic map has a unique asymptotically stable invariant measure.
Recall that the logistic family: $f_{\lambda}(x)=\lambda x(1-x)$ undergoes a period-doubling cascade as $\lambda$ increases from 1 to approximately 3.56995. For higher values of $\lambda$, the family exhibits stable periodic orbits of odd period, following the Sarkowski sequence, culminating with a stable period-3 orbit when $\lambda=1+\sqrt{8}$. When $\lambda$ is greater than this value the maps exhibit chaotic behavior on a Cantor set of parameter values. In this section, we are interested in randomly perturbing the values of $\lambda$ in this chaotic regime: $\lambda\in(1+\sqrt{8},4)$. The logistic family is well-studied in the literature, and more details on its dynamics can be found in, for example, [16]
In [17], Athreya and Dai consider the stochastic logistic map and demonstrate that the Markov process $\{X_{n}\}$ is Feller continuous, which guarantees the existence of the invariant distribution (but not its stability) by Theorem
1. They also show that if $E[\log\lambda_{1}]>0,E[\log(4-\lambda_{1})]<\infty$, and if the process is Harris irreducible with respect to some $sigma$-finite measure , then the probability distribution converges in the sense of Césaro to a unique probability distribution on (0,1). We shall consider the case when $\lambda\sim\mathcal{U}(3.87,4)$ to ensure that the maps are in the chaotic regime. We show that, in this case, the stochastic logistic map satisfies the conditions laid out in [17] for convergence of the distribution in the sense of Céaround.
Remark 1.
Let $\lambda\sim\mathcal{U}(3.87,4)$. Then we claim that $E[\log\lambda]>0$ and $E[\log(4-\lambda)]<\infty$. Notice that if $\lambda\sim\mathcal{U}(3.87,4)$, then $E(\lambda)=3.935>0$. Similarly,
[…]
$\approx$ $-12,$
Thus $E[\log(4-\lambda)]$ is finite.
This remark implies that if the process $\{X_{n}\}$ is Harris irreducible, then there exists a unique probability distribution on $(0,1)$ to which other probability distributions converge to in the sense of Cesaro.
The following results prove the Harris irreducibility and aperiodicity of the Markov process, which then gives the asymptotic stability of the invariant measure.
Lemma 1
and its generalization
Lemma 2 build toward Theorem 4, which gives the Harris irreducibility of the Markov process. Theorem 5 shows that the process is strongly aperiodic and, hence, has a unique asymptotically stable invariant measure.
Lemma 1.
Let $\lambda_{i}\sim\mathcal{U}(3.87,4)$ for all $i\geq 1$, $\alpha$ a fixed element in $(3.87,4)$, and $x^{*}$ such that $x^{*}=\alpha x^{*}(1-x^{*})$. (Equivalently, $x^{*}=1-\frac{1}{\alpha}$.) Then there exists $\eta>0$ such that for all $x\in I=(x^{*}-\eta,x^{*}+\eta)$ and $A\subset I$ with $m(A)>0$, then
$p(x,A)=\hbox{Prob}\{X_{1}\in A\,|\,X_{0}=x\}>0.$
where $m$ is Lebesgue measure.
That is, there is a neighborhood $I$ around the fixed point for which every $x\in I$ has a positive probability of visiting each subset of $I$ of positive
Lebesgue
measure.
Proof.
Fix $\alpha\in(3.87,4)$, and let $x^{*}\in(0,1)$ be such that $\alpha x^{*}(1-x^{*})=x^{*}$.
Notice that for any $\eta>0$ and $A\subset I\equiv(x^{*}-\eta,x^{*}+\eta)$ with $m(A)>0$
[…]
and
[…]
Since $h(x)=x(1-x)$ is continuous at $x^{*}$ and $\alpha h(x^{*})=x^{*}$, it follows that there exists $\eta^{*}>0$ such that $a(\eta^{*})<4$ and $b(\eta^{*})>3.87$.   Let $A_{\eta^{*}}\subset I_{\eta^{*}}=(x^{*}-\eta^{*},x^{*}+\eta^{*})$ be a set of positive Lebesgue measure associated with this particular value of $\eta$. Then, for any $x\in S$, ${\frac{A_{\eta^{*}}}{x(1-x)}}\subset(3.87,4)$, and since $A_{\eta^{*}}$ has positive measure, so does $\frac{A_{\eta^{*}}}{x(1-x)}$, and therefore $p(x,A_{\eta^{*}})>0$.
Remark 2.
Note that while we take the distribution of $\lambda$ on $(3.87,4)$ to be uniform, the same result will hold if we replace the uniform distribution with a distribution that has non-vanishing absolutely continuous part
on this interval.
The next lemma generalizes Lemma
1
to the case in which the map has a stable periodic orbit of any period.
Lemma 2.
As before, let $\lambda_{i}\sim\mathcal{U}(3.87,4)$ for all $i\geq 1$. Suppose $\alpha$ is a parameter value in $(3.87,4)$ with an $m$-periodic point, $p$, so that $f^{(m)}(p,\alpha)=p$. Then there is an $\eta>0$ such that, for any $x\in I=(p-\eta,p+\eta)$ and $A\subset I$ with $m(A)>0$, we have
$p^{m}(x,A)=Prob(X_{m}\in A|X_{0}=x)>0$
Proof.
Since $43.87<\alpha<$ 4 we can choose an $\eta_{1}$ small enough so that
$43.87<\left({\frac{p-\eta_{1}}{p+\alpha\eta_{1}}}\right)\alpha<\left({\frac{p+\eta_{1}}{p-\alpha\eta_{1}}}\right)\alpha<$ 4
because
$\lim_{\eta_{1}\rightarrow 0}\frac{p-\eta_{1}}{p+\alpha\eta_{1}}=\lim_{\eta_{1}\rightarrow 0}\frac{p+\eta_{1}}{p-\alpha\eta_{1}}=1$
Denote points in the periodic orbit of $p$ as $p_{j}=f^{(j)}(p,\alpha)$. Since $p$ has period $m$, $f(p_{m-1},\alpha)=\alpha p_{m-1}(1-p_{m-1})=p$. Since $f(\cdot,\cdot)$ is continuous in both variables:
  1. 1.
    there is an $\eta_{2}>0$ such that, if $|y-p_{m-1}|<\eta_{2}$ then
    $|f(y,\alpha)-p|<\eta_{1}\alpha\Leftrightarrow|y(1-y)-\frac{p}{\alpha}|<\eta_{1}\,;$
  2. 2.
    given $\eta_{2}$ as in the previous statement, $\exists\,\eta_{3}$ so that $|x-p|<\eta_{3}$ and $|\lambda_{i}-\alpha|<\eta_{3}$ for $1\leq i<m\Rightarrow|X_{m-1}-p_{m-1}|<\eta_{2}$ where $X_{j+1}=\lambda_{j+1}X_{j}(1-X_{j})$ and $X_{0}=x$.
In other words, if the random logistic map starts close to $p$, and if all of the parameter values are close to $\alpha$, then the random orbit returns close to $p$ after $m$ iterates.
Let $0<\eta<\min\{\eta_{1},\eta_{3}\}$. Then, if $|\lambda_{i}-\alpha|<\eta$ for $1\leq i<m$, and $A\subset I=(p-\eta,p+\eta)$, with $X_{0}\in I$, we have:
$\frac{A}{X_{m-1}(1-X_{m-1})}\subset\left({\frac{p-\eta_{1}}{\frac{p}{\alpha}+\eta_{1}},\frac{p+\eta_{1}}{\frac{p}{\alpha}-\eta_{1}}}\right)\subset(3.87,4)$
by our original assumption on $\eta_{1}$. Since $A$ has positive measure, so does ${\frac{A}{X_{m-1}(1-X_{m-1})}}$. To help clarify the notation, let $C$ be the event that $\lambda_{i}\in(\alpha-\eta,\alpha+\eta)$ for $1\leq i<m$. What we have shown is that
$Prob(X_{m}\in A|X_{m-1},C)=Prob\left({\lambda_{m}\in\frac{A}{X_{m-1}(1-X_{m-1})}|X_{m-1},C}\right)>0.$ (3)
Assume that $X_{0}=x\in I$, and recall that $\lambda_{m}$ is independent of $X_{0},X_{1},\dots X_{m-1}$. Then by the law of total probability:
$Prob(X_{m}\in A)\geq Prob((X_{m}\in A)\cap C)=E_{x}[Prob((X_{m}\in A)\cap C)|(X_{m-1}=x)]$ (4)
where $E_{x}$ is the expected value with respect to the distribution of $X_{m-1}$. Focusing on what’s inside the expected value:
$Prob\left({(X_{m}\in A)\cap C|X_{m-1}}\right)=Prob(X_{m}\in A|X_{m-1},C)Prob(C)$
Since the $\lambda$’s are independent of $X_{i}$ and $Prob(|\lambda_{i}-\alpha|<\eta)>0$, we have by Equation (3)
$Prob\left({(X_{m}\in A)\cap C|X_{m-1}}\right)=Prob(X_{m}\in A|X_{m-1},C)\left[{Prob(|\lambda_{i}-\alpha|<\eta)}\right]^{m-1}>0$
The expected value of a strictly positive random variable is positive, so we conclude, using Equation (4) that, if $X_{0}\in I$ then $Prob(X_{m}\in A)>0$, as claimed. ∎
Remark 3.
Lemma 2 shows that, if there is a parameter $\alpha$ in the support of the parameter values for which the map has a stable periodic orbit, then the process $\{X_{n}\}$ is Harris irreducible on $S$ with reference measure $\phi(\cdot)=m(\cdot\cap I)$, where $m$ is Lebesgue measure.
Theorem 4.
f As above, let $S=[0,1]$.
$X_{n+1}=\lambda_{n+1}X_{n}(1-X_{n}),\,\,\,n\geq 0$
with $\{\lambda_{n}\}_{n\geq 1}$ being i.i.d., $\lambda_{i}\sim\mathcal{U}(3.87,4)$, and $X_{0}$ an independent $S$-valued r.v. Then $\{X_{n}\}$ is irreducible with respect to the Lebesgue measure $m$ on $S$.
Proof.
By a result of Graczyk and Swiatek [18], the set of all $\lambda$ in $(0,4)$ such that $f(\cdot,\lambda)$ has a stable periodic orbit is dense. Thus, there is a $\alpha\in(3.87,4)$ and $10<x^{*}<$ 1 such that $x^{*}$ is a stable periodic point for $f(x,\alpha)=\alpha x(1-x)$. Let $\delta$ be such that $(\alpha-\delta,\alpha+\delta)\subset(3.87,4)$.
Let $K=\{x\in(0,1)\,\,|\,\,\omega_{\alpha}(x)=\gamma(x^{*})\}$, where $\omega_{\alpha}$ is the limit set of the orbit of $x$ under $f(\cdot,\alpha)$ and $\gamma(x^{*})$ is the orbit of the stable periodic point $x^{*}$. In [19] (Chapter V, Theorem 1.3) de Melo and van Strien have shown that $m(K)=1$ for unimodal maps, where $m$ is Lebesgue measure. Thus let $I=(x^{*}-\eta,x^{*}+\eta)$ as given by Lemma 2. Then, for all $x\in K$, $f^{(n)}(x,\alpha)\in I$ for some large $n$. Thus, by continuity of $f$,
$Prob(X_{n}\in I\text{ for some }n\geq 1|X_{0}=x)>0$ (5)
for all $x\in K$.
We will next show that for any $x\in(0,1)$, there is a positive probability that $x$ is mapped to $K$ in one iterate. In fact, we will show that $p(x,K)=Prob(X_{1}\in K|X_{0}=x)=1$. Fix $x\in(0,1)$ and let $K^{c}$ be the complement of $K$ in $[0,1]$. Since $\lambda$ is uniformly distributed on $[3.87,4]$ and $m(K^{c})=0$ we have
$p(x,K)$ $=$ $Prob(\lambda\,|\,\lambda x(1-x)\in K)$
$=$ $Prob\left({\lambda\in\frac{K}{x(1-x)}}\right)$
$=$ $Prob\left({(3.87,4)\cap\frac{K}{x(1-x)}}\right)$
$=$ $1-Prob\left({(3.87,4)\cap\frac{K^{c}}{x(1-x)}}\right)$
$\geq$ $1\displaystyle 1-Prob\left({\frac{K^{c}}{x(1-x)}}\right)=$ 1
Therefore, $p(x,K)>0$, establishing irreducibility. ∎
We now have that, since $E(\lambda)>0$ and $E(\log(4-\lambda))<\infty$, and the process is Harris irreducible, by [17], the process $\{X_{i}\}$ has a unique invariant measure that is stable in the sense of Cesaro. We conclude this section by demonstrating that the process is strongly aperiodic, and thus the invariant distribution is asymptotically stable with respect to the total variation norm.
Theorem 5.
Let $S=[0,1]$.
$X_{n+1}=\lambda_{n+1}X_{n}(1-X_{n}),\,\,\,n\geq 0$
with $\{\lambda_{n}\}_{n\geq 1}$ being i.i.d., $\lambda_{i}\sim\mathcal{U}(3.87,4)$, and $X_{0}$ an independent $S$-valued r.v. Then $\{X_{n}\}$ is strongly aperiodic, and thus has an asymptotically stable invariant distribution.
Proof.
Let $\lambda_{min}=3.87$, and let $A_{0}=(1-\frac{1}{\lambda_{min}},1-\frac{1}{4})$ (that is, $A_{0}$ is the set of fixed points for the deterministic logistic maps corresponding to the possible parameter $\lambda$ values). Let $L_{1}$ be the length of $A_{0}$ (that is, $L_{1}=\frac{1}{\lambda_{min}}-\frac{1}{4}$ ) and let $L_{2}$ be the length of interval $(\lambda_{min},4)=4-\lambda_{\min}$. Let $\nu$ be normalized Lebesgue measure on $A_{0}$. Then $A_{0}$, by definition, has measure 1, and for any measurable set $A$ :
[…]
Fix $x\in A_{0}$ and $A\subset S$. We want to show that $p(x,A)\geq c\,\nu(A)$ for some positive constant, $c$ (Definition 5). We have:
$p(x,A)$ $=$ $Prob\left(\lambda\in\frac{A}{x(1-x)}\right)$
[…]
$=$ $\frac{1}{L_{2}}\cdot\frac{1}{x(1-x)}\cdot L_{1}\nu(A)$
$\geq$ $4\cdot\frac{L_{1}}{L_{2}}\nu(A),$
since $x(1-x)\leq\frac{1}{4}$ for all $x\in[0,1]$. Thus, letting $\alpha=4\cdot\frac{L_{1}}{L_{2}}$ gives us the desired result. Therefore, the process $\{X_{n}\}$ is strongly aperiodic, and has an asymptotically stable invariant distribution. ∎
4 Simulations and Discussion
All code that was used to generate the figures in this section can be found at the GitHub repository here: https://github.com/kdayers/NoisyFixedPoints. Suppose $X$ is a random variable taking values in $[0,1]$. If $Pr\{X\leq x\}=Pr\{f(X)\leq x\}$ then we say that the distribution of $X$ is invariant under $f$. In this way, we can discuss invariant distributions for the deterministic logistic map, by looking at distribtuions invariant under $f_{\lambda}(x)=\lambda x(1-x)$. In [20], Ulam and von Neumann proved that the continuous invariant distribution for the deterministic logistic map with $\lambda=4$ is a beta distribution with $\alpha=\beta=0.5$. This function is symmetric on the interval (0,1), and goes to infinity at the endpoints of its support. Thus, the invariant distribution when $\lambda=4$ is given by the function
$f^{*}(x)=\begin{cases}\frac{1}{\pi\sqrt{x(1-x)}}&0<x<1\\ 0&\hbox{ else}\end{cases}$ (6)
See Figure 1 for a comparison of this beta density and an empirical distribution.
Figure 1: Histogram of the beta distribution with $\alpha=\beta=0.5$. The solid line shows the theoretical distribution given in Equation (6); the bars are a histogram generated by sampling 100,000 initial values uniformly on $(0,1)$
However, invariant densities for other values of $\lambda$ are not, in general, symmetric. In [21], Hall and Wolff provided numerical estimates for the invariant distributions for the deterministic logistic map for various values of $\lambda$. They demonstrated that any continuous invariant density will have poles (singularities) at the boundary of its support, as well as additional poles at the interior of the support in the case $\lambda<4$. They also show that these invariant distributions were not symmetric. Figure 2 shows numerical approximations to theses deterministic invariant densities for $\lambda=3.87$ and $\lambda=3.95$ These numerical approximations are histograms of 100,000 sampled initial values iterated 2,000 times. They indicate the singularities in the theoretical densities described in Remark 5 of [21]. These singularities, or poles, correspond to the orbit of the critical point, $f_{\lambda}(1/2)$. In addition, Hall and Wolff observed that the invariant density will be supported on the interval $(\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}),\frac{1}{4}\lambda)$. This is due to the fact that any point in the interval $(\frac{1}{4}\lambda,1)$ will be mapped into the interval $(0,\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}))$, and any point in $(0,\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4})$ will be mapped to a point in $[\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}),\frac{1}{4}\lambda]$. Therefore, after finitely many iterations, any point $x\in(0,1)$ must be mapped into $[\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}),\frac{1}{4}\lambda]$, and so the invariant density must be supported on this subinterval. Note that as $\lambda\rightarrow 4$, $\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4})\rightarrow 0$ and $\frac{1}{4}\lambda\rightarrow 1$, so the interval of support widens as $\lambda$ approaches 4.
Figure 2: Histogram of the simulated invariant distribution for the deterministic logistic map with $\lambda=3.87$ (left) and $\lambda=3.935$ (right) for 100,000 sampled values, after $n=20$ iterates.
To numerically explore convergence to the invariant distribution in the stochastic case, we ran several simulations with 100,000 initial values randomly sampled according to a variety of distributions:
  1. 1.
    A discrete distribution supported on the set $\{0.11,0.33,0.55,0.6,0.78\}$ with probabilities $\{0.196,0.140,0.233,0.322,0.107\}$ respectively (Figure 3, left).
  2. 2.
    A discrete distribution supported on $\{0.25,0.5,0.75\}$ each with probability 1/3.
  3. 3.
    A truncated exponential distribution on (0,1) with parameter $r=1.2$ (Figure 3, right).
  4. 4.
    A truncated gamma distribution on (0,1) with shape parameter 3 and scale parameter 1.
  5. 5.
    A truncated normal distribution on (0,1) with mean 0.5 and standard deviation 0.3.
  6. 6.
    A truncated $t$-distribution on (0,1) with 1 degree of freedom.
  7. 7.
    A uniform distribution on (0,1) (Figure 3, center).
In each case, after 2,000 iterations, we obtained a histogram very similar to that presented in Figure 4: a bimodal, asymmetric distribution that appears to be supported throughout the unit interval, without singularities (poles) in the interior. This is noteworthy, particularly because we simulated using both continuous and discrete starting distributions. We examined distributions that were left-skewed, right-skewed, and symmetric. Each simulated initial distribution appears to converge to the same distribution after 2,000 iterations, aligning with the global asymptotic stability of the invariant distribution established theoretically earlier in this paper. We also note that, while the distribution demonstrates peaks toward the sides of the interval, it does not exhibit the poles at the boundary of the interval in the same manner as the invariant distributions for the deterministic maps. However, this seems reasonable, as the interval of support of the invariant distribution is dependent on $\lambda$, as discussed above, and relatively fewer values of $\lambda$ will map points into ranges closer to 0 and 1, so it makes sense that the distribution will decrease near the boundary of $[0,1]$. It is also of note that even as the width of the interval that the distribution of $\lambda$ is supported on narrows, the invariant distribution maintains this property of decreasing near the boundary of support. For instance, we ran simulations where $\lambda$ was sampled uniformly on the interval $(3.87,9)$ and again on the interval $(3.87,3.935).$ Both invariant distributions exhibited this decrease toward the edge of the boundary of support (which, as expected, is not the entirety of $(0,1)$. Please see Figure 5.
As for the speed of convergence, it appears that the convergence happens relatively quickly. Figure 6 shows the evolution of the simulated invariant density after various numbers of iterations, starting from a uniform distribution.
Refer to caption Refer to caption Refer to caption
Figure 3: Starting distributions of 100,000 starting values sampled on the set $\{0.11,0.33,0.55,0.6,0.78\}$ with probabilities $\{0.196,0.140,0.233,0.322,0.107\}$ respectively (left), the uniform distribution on (0,1) (middle), and the truncated exponential distribution with parameter $r=1.25$ (right).
Figure 4: Empirical distribution of $\{X_{n}\}$ after $n=2,000$ iterations, where 100,000 initial values were sampled from a truncated exponential distribution with parameter $r=1.25$ as in Figure 3. Similar results were obtained using other initial distributions (results not shown).
Figure 5: The simulated invariant distribution of $\{X_{n}\}$ for $\lambda\sim U(3.87,3.9)$ (left) and $\lambda\sim U(3.87,3.935)$ (right), shown after 20 iterations.
Figure 6: The densities for a simulation of 100,000 values, initially sampled randomly from a uniform distribution on (0,1), after various numbers of iterations.
5 Conclusion
In conclusion, we have illustrated that the stochastic logistic map, characterized by parameter values uniformly distributed in the chaotic regime, exhibits Harris irreducibility. Consequently, the previously established invariant distribution stands as the unique invariant distribution. This result also implies that all initial distributions converge to the invariant distribution in the Cesàro sense. Moreover, we have demonstrated the strong aperiodicity of the stochastic logistic map, indicating that initial distributions converge to the invariant distribution with respect to the total variation distance—a more robust and rigorous result. We conducted simulations of the stochastic logistic map to visualize the invariant distribution. Remarkably, the empirically derived invariant distribution exhibits qualitative distinctions when compared to the invariant distributions of the deterministic logistic map for values of $\lambda$ in the chaotic regime. In particular, the simulated distribution does not appear to have a pole at $x=0$, which contrasts with the work in [21], which demonstrates that the deterministic invariant density will have a pole at the boundary of its support.
It should be noted that this work does not seem to depend on the $\lambda$ values being uniformly distributed on (3.87,4) and indeed, it appears that the results will hold for any absolutely continuous distribution on (3.87,4).
In future work, we plan to provide a theoretical basis for these empirical observations. In particular, we hope to better describe the asymptotically stable invariant distribution in the case that the $\lambda$-values are uniformly distributed (the case illustrated in Figure 4). We also plan to provide similar theoretical results when the parameter has other distributions, such as an exponential or discrete distribution.

### v3

@@
 A natural question is: given an initial value, $x_{0}$, what is the long-term behavior of the sequence $x_{0},x_{1},x_{2},\dots$?
-If the sequence converges to a finite value, it must be a fixed point, i.e.
-a value $x^{*}$ that satisfies $x^{*}=f(x^{*})$.
+If the function $f$ is continuous, and the sequence converges to a finite value, $x^{*}$, then it must be a fixed point, i.e.
+$x^{*}=f(x^{*})$.
 So one of the first questions a dynamicist asks when encountering a dynamical system is: where are the fixed points?
@@
 If such a distribution exists, we are also interested in its stability; that is, whether other distributions converge to this invariant distribution in some manner.
-We define the random logistic map and the notions of invariant measures and stability more precisely below.
+We define the stochastic logistic map and the notions of invariant measures and their stability more precisely below.
 In this paper, we consider the state space $S=[0,1]$ and the particular choice of parameter space $\Lambda=(3.87,4)\subset[0,4]$.
@@
 This will be important in the proofs in Section 3.
-Let $\mathcal{B}(S)$ denote the Borel sigma-algebra on $S$. We consider the logistic Markov process
+Let $\mathcal{B}(S)$ denote the Borel $\sigma$-algebra on $S$. We consider the logistic Markov process
 $X_{n+1}=\lambda_{n+1}X_{n}(1-X_{n}),$ (1)
@@
 $\hbox{Prob}\{X_{n}\in A|X_{0}=x\}=p^{n}(x,A).$
-Suppose we have a distribution, $\mu$, on the state space, $S$.
+Suppose we have a distribution, $\mu$, on the state space, $S$, which gives the distribution of $X_{n}$.
 When we apply the process (1) to subsets of $S$, we induce a new distribution, $P\mu$ on the state space as follows.
@@
 We can iterate the operator $P$ a positive number $m$ times, which we denote by $P^{m}$: if $X_{0}$ has distribution $\mu$ then $X_{m}$ has distribution $P^{m}\mu$.
-This leads us the analog to stability in the stochastic context.
+This leads us to the analog of stability in the stochastic context.
 Definition 2.
@@
 […]
-In the deterministic case, we know that, if the iterates of $x_{0}$ under a continuous function, $f$, converge to a point, $x$, then $x$ must be a fixed point of $f$.
-This result follows from the continuity of $f$.
+In the deterministic case, we know that, if the iterates of $x_{0}$ under a continuous function, $f$, converge to a point, $x$, then $x$ must be a fixed point of $f$, and this result follows from the continuity of $f$.
 We can get a similar result in the stochastic setting, but we first need to define a notion of continuity.
@@
 is continuous.
-Using weak continuity we get the following result ([14], Theorem 11.1).
+Using weak continuity we get the following result ([14], Theorem 11.1 of Chapter 2).
 Proposition 1.
@@
     If there is an $x\in S$, a sequence of integers $1\leq n_{1}<n_{2}<\dots$ and a probability measure $\pi_{x}$ on $S$ such that
-    $\frac{1}{n_{k}}\sum_{m=0}^{n_{k}-1}p^{(m)}(x,dy)\hbox{ converges weakly to }\pi_{x}\hbox{ as }k\rightarrow\infty$
+    $\frac{1}{n_{k}}\sum_{m=0}^{n_{k}-1}p^{m}(x,dy)\hbox{ converges weakly to }\pi_{x}\hbox{ as }k\rightarrow\infty$
     then $\pi_{x}$ is an invariant measure for $p$.
   2. 2.
-    If the limit above exists for all $x\in S$ and is independent of $x$ then $\pi=\pi_{x}$ is the unique invariant probability for $p$.
-We can state this in terms of the operator $P$ given in Equation (2).
+    If, for all $x\in S$, there is a sequence of integers $\{n_{i}\}$ such that the limit above exists and is independent of $x$ then $\pi=\pi_{x}$ is the unique invariant probability for $p$.
+We can state this proposition in terms of the operator $P$ given in Equation (2).
 Suppose a process has the Feller property and we have a measure $\mu$ on $S$.
@@
 $\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{m=0}^{n-1}P^{m}\mu(A){\rightarrow}\pi(A)$
-then $\pi$ is an invariant measure under the operator $P$, i.e. $P\pi(A)=\pi(A)$ for $A\in{\cal B}(S)$. Furthermore, if Condition 2 holds in Proposition 1, then this invariant measure is unique.
+then $\pi$ is an invariant measure under the operator $P$, i.e.
+$P\pi(A)=\pi(A)$ for $A\in{\cal B}(S)$.
+Furthermore, if the condition in Item 2 of Proposition 1 holds, then this invariant measure is unique.
 What we would like is the asymptotic stability of the (unique) invariant measure, $\pi$. In other words, for any probability measure $\mu$ on $S$, when is it the case that
@@
 Let $\mathcal{B}(S)$ denote the Borel subsets of $S$.
-A Markov process $\{X_{n}\}_{n=0}^{\infty}$ is Harris irreducible with reference measure $\nu$ (or $\nu$-irreducible for short) if, for every $x\in S$ and every $B\in\mathcal{B}(S)$ with $\nu(B)>0$, there is $n\geq 1$ such that $p^{n}(x,B)>0$.
-In words: a Markov process is Harris irreducible if every element in the state space eventually gets mapped to any set of positive probability.
-In the theorems that follow, we require that the process be Harris irreducible with respect to any $sigma$-finite measure: the particular measure is not important.
-Therefore, to make the language more concise, we sometimes just say “Harris irreducible” to mean “Harris irreducible with respect to some $\sigma$-finite measure.”
+  1. (i)
+    A Markov process $\{X_{n}\}_{n=0}^{\infty}$ is irreducible with respect to a non-zero $\sigma$-finite reference measure $\nu$ (or $\nu$-irreducible for short) if, for every $x\in S$ and every $B\in\mathcal{B}(S)$ with $\nu(B)>0$, there is $n\geq 1$ such that $p^{n}(x,B)>0$.
+  2. (ii)
+    The Markov process is $\phi$-recurrent if, for each $x\in S$ and $A\in\mathcal{B}(S)$ such that $\phi(A)>0$, we have
+    $P(\eta_{A}<\infty|X_{0}=x)=1,\quad\hbox{ where }\;\eta_{A}=\inf\{n\geq 1:X_{n}\in A\}$
+  3. (iii)
+    If the Markov process is both $\nu$-irreducible and $\phi$-recurrent, then it is Harris recurrent.
+  4. (iv)
+    If a Harris recurrent process satisfies
+    $\sup\{E(\eta_{A_{0}}|X_{0}=x):x\in A_{0}\}<\infty$
+    for some $\phi$-recurrent set $A_{0}$, then the process is said to be positive Harris recurrent.
+Note: the set $A_{0}$ is $\phi$-recurrent if $\phi(A_{0})>0$ and it is recurrent, i.e. $P(\eta_{A_{0}}<\infty|X_{0}=x)=1$ $\forall x\in{\mathcal{B}}(S)$.
+In other words: a Markov process is $\nu$-irreducible if there is a positive probability that every element in the state space eventually gets mapped to any set of positive $\nu$ measure.
+It is $\phi$-recurrent if, with probability 1, it visits every set of positive $\phi$-measure in a finite number of steps.
 The Krylov-Bogolubov Theorem [3] tells us that a stochastic dynamical system obtained from a continuous function on a compact subset of ${\mathbb{R}}^{d}$ with independent parametric noise always has an invariant distribution:
@@
 A Feller continuous, stationary Markov process on a compact subset of ${\mathbb{R}}^{d}$ has an invariant distribution.
-The theorem applies to more general settings, but this more specific version is all we need here. Harris irreducibility can then be used to show that this invariant distribution is unique.
-Theorem 2 ([14], Theorem 2.C9.3).
-If a Markov process is Harris irreducible with respect to some $\sigma$-finite measure $\nu$ on $S$ and has an invariant probability $\mu^{*}$, then $\mu^{*}$ is the unique invariant probability and, for any probability measure $\mu$ on S:
+The theorem applies to more general settings, but this more specific version is all we need here.
+Irreducibility can then be used to show that this invariant distribution is unique using the following theorem:
+Theorem 2.
+[[14], Theorem C9.3 in Chapter 2] If a Markov process is irreducible with respect to some $\sigma$-finite measure $\nu$ on $S$ and has an invariant probability $\mu^{*}$, then the process is positive Harris recurrent, $\mu^{*}$ is the unique invariant probability and, for any probability measure $\mu$ on S:
 $\lim_{n\rightarrow\infty}\sup_{A\in\mathcal{B}(S)}\left|\frac{1}{n}\sum_{m=1}^{n}P^{m}\mu(A)-\mu^{*}(A)\right|=0$
-Thus, if a Markov process is Harris irreducible, we know that any measure converges to a unique invariant measure, $\mu^{*}$, under the operator, $P$ “in the sense of Césaro”, i.e.
-the average of the sequence of measures conerges to $\mu^{*}$.
-To get asymptotic stability of the invariant meausure without this average, we need one more condition: aperiodicity.
+Thus, if a Markov process is irreducible and has some invariant measure, we know that any measure converges to a unique invariant measure, $\mu^{*}$, under the operator, $P$ “in the sense of Césaro”, i.e.
+the average of the sequence of measures converges to $\mu^{*}$.
+We also get positive Harris recurrence from this result.
+To get asymptotic stability of the invariant measure without this average, we need one more condition: aperiodicity.
 Definition 5.
-A Markov process is said to be strongly aperiodic if there is a set $A_{0}\subset S$ with $\nu(A_{0})=1$ and an $\alpha>0$ such that for all $x\in A_{0}$, $A$ a measurable subset of $S$, one has
-$p(x,A)\geq\alpha\nu(A)$
-Theorem 3 ([14], Theorem C9.2).
-If a Markov process is Harris irreducible with respect to some $\sigma$-finite measure $\nu$ on $S$ and strongly aperiodic with invariant measure $\mu^{*}$ then, for any $\mu\in\mathcal{B}(S)$,
-$\lim_{n\rightarrow\infty}\delta(P^{n}\mu,\mu^{*})=0$
+A Markov process is said to be strongly aperiodic with respect to a $\sigma$-finite probability measure $\phi$ on $S$ if there is a set $A_{0}\subset S$ with $\phi(A_{0})=1$ and an $\alpha>0$ such that one has
+  1. 1.
+    for all $x\in S$, $P(\eta_{A_{0}}<\infty|X_{0}=x)=1$, i.e. the process is $A_{0}$-recurrent;
+  2. 2.
+    for all $x\in A_{0}$, $A$ a measurable subset of $S$, $p(x,A)\geq\alpha\phi(A)$.
+Let $\{X_{n}\}$ be a Markov process on a state space $S$, and ${\mathcal{B}}$ the Borel subsets of $S$.
+Theorem 3 ([15], Corollary 6.7).
+Suppose $\{X_{n}\}$ is a positive Harris recurrent and aperiodic Markov process with stationary distribution $\pi$. Then for any initial distribution, $\mu$, of $X_{0}$:
+$\lim_{n\rightarrow\infty}\delta(P^{n}\mu,\pi)=0$
 where $\delta$ is the total variation distance
 $\delta(P,Q)=\sup_{B\in\mathcal{B}}|P(B)-Q(B)|.$
+We note that the measures used in the definitions of irreducibility and aperiodicity could be different.
+(This fact will be used later in the proof of Theorem 5.) To see that this is the case, consider the following.
+Aperiodicity with respect to the measure $\nu$ implies that there is an $\alpha>0$ such that
+$p(x,B)\geq\alpha\nu(B)\,\Rightarrow\,p(x,B)>0\;\hbox{ if }\nu(B)>0$
+Thus, the process is $\nu$-irreducible. Therefore, in the statement of Theorem 3, we do not need to specify the measures used to show either irreducibility or aperiodicity.
 As an example of the ideas presented in this section, let us consider a version of the stochastic logistic map as examined in [16].
@@
 If $1<\alpha<\beta\leq 2$ then there is a unique invariant probability on (0,1), and this probability is nonatomic.
-If $2<\alpha<\beta<1+\sqrt{5}$ and $\alpha\in[8/\beta(4-\beta),\beta)$, then there exists a unique invariant probability on (0,1) which is nonatomic and has its support contained in $[1/2,\lambda/4]$.
+If $2<\alpha<\beta<1+\sqrt{5}$ and $\alpha\in[8/\beta(4-\beta),\beta)$, then there exists a unique invariant probability on (0,1) which is nonatomic and has its support contained in $[1/2,(1+\sqrt{5})/4]$.
 3 Proof of the asymptotic stability of the invariant measure in the chaotic regime.
@@
 In this section we use these results to prove that, when parameters are supported in the chaotic regime, the random logistic map has a unique asymptotically stable invariant measure.
-Recall that the logistic family: $f_{\lambda}(x)=\lambda x(1-x)$ undergoes a period-doubling cascade as $\lambda$ increases from 1 to approximately 3.56995.
-For higher values of $\lambda$, the family exhibits stable periodic orbits of odd period, following the Sarkowski sequence, culminating with a stable period-3 orbit when $\lambda=1+\sqrt{8}$.
-When $\lambda$ is greater than this value the maps exhibit chaotic behavior on a Cantor set of parameter values.
-In this section, we are interested in randomly perturbing the values of $\lambda$ in this chaotic regime: $\lambda\in(1+\sqrt{8},4)$.
+Recall that the deterministic logistic family: $f(x,\lambda)=f_{\lambda}(x)=\lambda x(1-x)$ undergoes a period-doubling cascade as $\lambda$ increases from 1 to $\lambda_{2\infty}\approx$ 3.56995.
+Note that here iteration occurs only with respect to $x$; that is $f^{(m)}(x,\lambda)=f(f^{m-1}(x,\lambda),\lambda)$, where $\lambda$ remains fixed.
+For some higher values of $\lambda$, the family exhibits stable periodic orbits of odd period, following the Sarkowski sequence, culminating with a stable period-3 orbit when $\lambda=1+\sqrt{8}$.
+Interspersed between the values of $\lambda$ with stable periodic orbits are a set of large measure of parameter values for which there are no stable periodic orbits, and almost all initial conditions give rise to orbits with no finite period.
+Thus, $\lambda_{2\infty}$ marks the “onset of chaos”.
+As $\lambda$ increases beyond $1+\sqrt{8}$, the family of maps again experiences a period-doubling cascade, with stable orbits of period 6, 12, 24, 48, …, culminating at $\lambda_{3\infty}\approx 3.8495$.
+When $\lambda$ is greater than this value, the maps exhibit chaotic behavior on a Cantor set of parameter values, with three “chaotic bands” of orbits merging into one chaotic band at $\lambda\approx 3.857$.
+In what follows, we are interested in randomly perturbing the values of $\lambda$ in this chaotic regime.
+To make sure we are well within the parameter range for which there is generically a single band chaotic attractor, we restrict $\lambda\in(3.87,4)$.
 The logistic family is well-studied in the literature, and more details on its dynamics can be found in, for example, [17]
-In [17], Athreya and Dai consider the stochastic logistic map and demonstrate that the Markov process $\{X_{n}\}$ is Feller continuous, which guarantees the existence of the invariant distribution (but not its stability) by Theorem
-1.
-They also show that if $E[\log\lambda_{1}]>0,E[\log(4-\lambda_{1})]<\infty$, and if the process is Harris irreducible with respect to some $sigma$-finite measure , then the probability distribution converges in the sense of Césaro to a unique probability distribution on (0,1).
-We shall consider the case when $\lambda\sim\mathcal{U}(3.87,4)$ to ensure that the maps are in the chaotic regime.
-We show that, in this case, the stochastic logistic map satisfies the conditions laid out in [17] for convergence of the distribution in the sense of Céaround.
-Remark 1.
-Let $\lambda\sim\mathcal{U}(3.87,4)$. Then we claim that $E[\log\lambda]>0$ and $E[\log(4-\lambda)]<\infty$. Notice that if $\lambda\sim\mathcal{U}(3.87,4)$, then $E(\lambda)=3.935>0$. Similarly,
-[…]
-$\approx$ $-12,$
-Thus $E[\log(4-\lambda)]$ is finite.
-This remark implies that if the process $\{X_{n}\}$ is Harris irreducible, then there exists a unique probability distribution on $(0,1)$ to which other probability distributions converge to in the sense of Cesaro.
-The following results prove the Harris irreducibility and aperiodicity of the Markov process, which then gives the asymptotic stability of the invariant measure.
-Lemma 1
-and its generalization
-Lemma 2 build toward Theorem 4, which gives the Harris irreducibility of the Markov process.
-Theorem 5 shows that the process is strongly aperiodic and, hence, has a unique asymptotically stable invariant measure.
+In [18], Athreya and Dai consider the stochastic logistic map and demonstrate that the Markov process $\{X_{n}\}$ is Feller continuous, which guarantees the existence of an invariant distribution (but not its stability) by Theorem 1.
+They also give conditions for convergence of the distribution of the $X_{i}$ in the sense of Césaro which differ somewhat from the hypotheses of Theorem 2.
+Here we choose to follow Bhattacharya and Majumbdar ([14]) and Nummelin ([15]) by showing the irreducibility and aperiodicity of the Markov process in the chaotic regime.
+This, along with the existence of an invariant measure, gives the asymptotic stability of the invariant measure by Theorem 3.
+Lemma 1 and its generalization Lemma 2 build toward Theorem 4, which shows that the Markov process is irreducible.
+Theorem 5 shows that the process is strongly aperiodic and thus, since it has an invariant measure by Theorem 1, the process has a unique asymptotically stable invariant measure.
 Lemma 1.
-Let $\lambda_{i}\sim\mathcal{U}(3.87,4)$ for all $i\geq 1$, $\alpha$ a fixed element in $(3.87,4)$, and $x^{*}$ such that $x^{*}=\alpha x^{*}(1-x^{*})$.
+Let $\lambda\sim\mathcal{U}(3.87,4)$, $\alpha$ a fixed element in $(3.87,4)$, and $x^{*}\in(0,1)$ such that $x^{*}=\alpha x^{*}(1-x^{*})$.
 (Equivalently, $x^{*}=1-\frac{1}{\alpha}$.) Then there exists $\eta>0$ such that for all $x\in I=(x^{*}-\eta,x^{*}+\eta)$ and $A\subset I$ with $m(A)>0$, then
@@
 where $m$ is Lebesgue measure.
-That is, there is a neighborhood $I$ around the fixed point for which every $x\in I$ has a positive probability of visiting each subset of $I$ of positive
-Lebesgue
-measure.
+That is, there is a neighborhood $I$ around the fixed point for which every $x\in I$ has a positive probability of visiting each subset of $I$ of positive Lebesgue measure.
 Proof.
@@
 Remark 1.
-Note that while we take the distribution of $\lambda$ on $(3.87,4)$ to be uniform, the same result will hold if we replace the uniform distribution with a distribution that has non-vanishing absolutely continuous part
-on this interval.
-The next lemma generalizes Lemma
-1
-to the case in which the map has a stable periodic orbit of any period.
+Note that while we take the distribution of $\lambda$ on $(3.87,4)$ to be uniform, the same result will hold if we replace the uniform distribution with a distribution that has non-vanishing absolutely continuous part on this interval.
+The next lemma generalizes Lemma 1 to the case in which the map has a stable periodic orbit of any period.
 Lemma 2.
 As before, let $\lambda_{i}\sim\mathcal{U}(3.87,4)$ for all $i\geq 1$.
-Suppose $\alpha$ is a parameter value in $(3.87,4)$ with an $m$-periodic point, $p$, so that $f^{(m)}(p,\alpha)=p$.
-Then there is an $\eta>0$ such that, for any $x\in I=(p-\eta,p+\eta)$ and $A\subset I$ with $m(A)>0$, we have
+Suppose $\alpha$ is a parameter value in $(3.87,4)$ with an $M$-periodic point, $x^{*}$, so that $f^{(M)}(x^{*},\alpha)=x^{*}$.
+Then there is an $\eta>0$ such that, for any $x\in I=(x^{*}-\eta,x^{*}+\eta)$ and $A\subset I$ with $m(A)>0$, we have
 $p^{M}(x,A)=Prob(X_{M}\in A|X_{0}=x)>0$
@@
 Since $43.87<\alpha<$ 4 we can choose an $\eta_{1}$ small enough so that
-$43.87<\left({\frac{p-\eta_{1}}{p+\alpha\eta_{1}}}\right)\alpha<\left({\frac{p+\eta_{1}}{p-\alpha\eta_{1}}}\right)\alpha<$ 4
+$43.87<\left({\frac{x^{*}-\eta_{1}}{x^{*}+\alpha\eta_{1}}}\right)\alpha<\left({\frac{x^{*}+\eta_{1}}{x^{*}-\alpha\eta_{1}}}\right)\alpha<$ 4
 because
-$\lim_{\eta_{1}\rightarrow 0}\frac{p-\eta_{1}}{p+\alpha\eta_{1}}=\lim_{\eta_{1}\rightarrow 0}\frac{p+\eta_{1}}{p-\alpha\eta_{1}}=1$
-Denote points in the periodic orbit of $p$ as $p_{j}=f^{(j)}(p,\alpha)$.
-Since $p$ has period $m$, $f(p_{m-1},\alpha)=\alpha p_{m-1}(1-p_{m-1})=p$.
+$\lim_{\eta_{1}\rightarrow 0}\frac{x^{*}-\eta_{1}}{x^{*}+\alpha\eta_{1}}=\lim_{\eta_{1}\rightarrow 0}\frac{x^{*}+\eta_{1}}{x^{*}-\alpha\eta_{1}}=1$
+Denote points in the periodic orbit of $x^{*}$ as $x^{*}_{j}=f^{(j)}(x^{*},\alpha)$.
+Since $x^{*}$ has period $M$, $f(x^{*}_{M-1},\alpha)=\alpha x^{*}_{M-1}(1-x^{*}_{M-1})=x^{*}$.
 Since $f(\cdot,\cdot)$ is continuous in both variables:
   1. 1.
-    there is an $\eta_{2}>0$ such that, if $|y-p_{m-1}|<\eta_{2}$ then
-    $|f(y,\alpha)-p|<\eta_{1}\alpha\Leftrightarrow|y(1-y)-\frac{p}{\alpha}|<\eta_{1}\,;$
+    there is an $\eta_{2}>0$ such that, if $|y-x^{*}_{M-1}|<\eta_{2}$ then
+    $|f(y,\alpha)-x^{*}|<\eta_{1}\alpha\Leftrightarrow\left|y(1-y)-\frac{x^{*}}{\alpha}\right|<\eta_{1}\,;$
   2. 2.
-    given $\eta_{2}$ as in the previous statement, $\exists\,\eta_{3}$ so that $|x-p|<\eta_{3}$ and $|\lambda_{i}-\alpha|<\eta_{3}$ for $1\leq i<m\Rightarrow|X_{m-1}-p_{m-1}|<\eta_{2}$ where $X_{j+1}=\lambda_{j+1}X_{j}(1-X_{j})$ and $X_{0}=x$.
-In other words, if the random logistic map starts close to $p$, and if all of the parameter values are close to $\alpha$, then the random orbit returns close to $p$ after $m$ iterates.
+    given $\eta_{2}$ as in the previous statement, $\exists\,\eta_{3}$ so that $|x-x^{*}|<\eta_{3}$ and $|\lambda_{i}-\alpha|<\eta_{3}$ for $1\leq i<M\Rightarrow|X_{M-1}-x^{*}_{M-1}|<\eta_{2}$ where $X_{j+1}=\lambda_{j+1}X_{j}(1-X_{j})$ and $X_{0}=x$.
+In other words, if the random logistic map starts close to $x^{*}$, and if all of the parameter values are close to $\alpha$, then the random orbit returns close to $x^{*}$ after $M$ iterates.
 Let $0<\eta<\min\{\eta_{1},\eta_{3}\}$. Then, if $|\lambda_{i}-\alpha|<\eta$ for $1\leq i<M$, and $A\subset I=(p-\eta,p+\eta)$, with $X_{0}\in I$, we have:
@@
 The expected value of a strictly positive random variable is positive, so we conclude, using Equation (4) that, if $X_{0}\in I$ then $Prob(X_{M}\in A)>0$, as claimed. ∎
-Remark 3.
-Lemma 2 shows that, if there is a parameter $\alpha$ in the support of the parameter values for which the map has a stable periodic orbit, then the process $\{X_{n}\}$ is Harris irreducible on $S$ with reference measure $\phi(\cdot)=m(\cdot\cap I)$, where $m$ is Lebesgue measure.
 Theorem 4.
-f As above, let $S=[0,1]$.
+As above, let $S=[0,1]$.
 $X_{n+1}=\lambda_{n+1}X_{n}(1-X_{n}),\,\,\,n\geq 0$
 with $\{\lambda_{n}\}_{n\geq 1}$ being i.i.d., $\lambda_{i}\sim\mathcal{U}(3.87,4)$, and $X_{0}$ an independent $S$-valued r.v.
-Then $\{X_{n}\}$ is irreducible with respect to the Lebesgue measure $m$ on $S$.
+Then $\{X_{n}\}$ is irreducible with respect to the Lebesgue measure $m$ on $S$ conditioned on $(x^{*}-\eta,x^{*}+\eta)$.
 Proof.
 By a result of Graczyk and Swiatek [19], the set of all $\lambda$ in $(0,4)$ such that $f(\cdot,\lambda)$ has a stable periodic orbit is dense.
-Thus, there is a $\alpha\in(3.87,4)$ and $10<x^{*}<$ 1 such that $x^{*}$ is a stable periodic point for $f(x,\alpha)=\alpha x(1-x)$.
-Let $\delta$ be such that $(\alpha-\delta,\alpha+\delta)\subset(3.87,4)$.
+Thus, there is an $\alpha\in(3.87,4)$ and $10<x^{*}<$ 1 such that $x^{*}$ is a stable periodic point for $f(x,\alpha)=\alpha x(1-x)$.
 Let $K=\{x\in(0,1)\,\,|\,\,\omega_{\alpha}(x)=\gamma(x^{*})\}$, where $\omega_{\alpha}$ is the limit set of the orbit of $x$ under $f(\cdot,\alpha)$ and $\gamma(x^{*})$ is the orbit of the stable periodic point $x^{*}$.
@@
 Fix $x\in(0,1)$ and let $K^{c}$ be the complement of $K$ in $[0,1]$.
-Since $\lambda$ is uniformly distributed on $[3.87,4]$ and $m(K^{c})=0$ we have
+Since $\lambda$ is uniformly distributed on $(3.87,4)$ and $m(K^{c})=0$ we have
 $p(x,K)$ $=$ $Prob(\lambda\,|\,\lambda x(1-x)\in K)$
@@
 Therefore, $p(x,K)>0$, establishing irreducibility. ∎
-We now have that, since $E(\lambda)>0$ and $E(\log(4-\lambda))<\infty$, and the process is Harris irreducible, by [17], the process $\{X_{i}\}$ has a unique invariant measure that is stable in the sense of Cesaro.
+Using Theorems 1 and 2, we now have that, since the process is Feller continuous by [18], the process $\{X_{i}\}$ has a unique invariant measure that is stable in the sense of Césaro.
 We conclude this section by demonstrating that the process is strongly aperiodic, and thus the invariant distribution is asymptotically stable with respect to the total variation norm.
@@
 Let $\lambda_{min}=3.87$, and let $A_{0}=(1-\frac{1}{\lambda_{min}},1-\frac{1}{4})$ (that is, $A_{0}$ is the set of fixed points for the deterministic logistic maps corresponding to the possible parameter $\lambda$ values).
-Let $L_{1}$ be the length of $A_{0}$ (that is, $L_{1}=\frac{1}{\lambda_{min}}-\frac{1}{4}$ ) and let $L_{2}$ be the length of interval $(\lambda_{min},4)=4-\lambda_{\min}$.
+Let $L_{1}$ be the length of $A_{0}$ (that is, $L_{1}=\frac{1}{\lambda_{min}}-\frac{1}{4}$ ) and let $L_{2}$ be the length of interval $(\lambda_{min},4)$ (that is, $L_{2}=4-\lambda_{\min}$).
 Let $\nu$ be normalized Lebesgue measure on $A_{0}$.
@@
 […]
-Fix $x\in A_{0}$ and $A\subset S$. We want to show that $p(x,A)\geq c\,\nu(A)$ for some positive constant, $c$ (Definition 5). We have:
+Fix $x\in A_{0}$ and $A\subset S$.
+We want to show that $p(x,A)\geq c\,\nu(A)$ for some positive constant, $c$ (Definition 5).
+Since $\nu(A)=\nu(A\cap A_{0})$ it suffices to demonstrate the above result for sets $A\subset A_{0}$.
+Notice for any $A\subset A_{0}$ and with $\nu(A)>0$, and $x\in A_{0}$,
+[…]
+and
+[…]
+Therefore, for any $A\subset A_{0}$ and $x\in A_{0}$, $\frac{A}{x(1-x)}\subset(3.87,4).$
+Given $x\in A_{0}$ and $A\subset A_{0}$, we therefore have:
 $p(x,A)$ $=$ $Prob\left(\lambda\in\frac{A}{x(1-x)}\right)$
@@
 Suppose $X$ is a random variable taking values in $[0,1]$.
-If $Pr\{X\leq x\}=Pr\{f(X)\leq x\}$ then we say that the distribution of $X$ is invariant under $f$.
-In this way, we can discuss invariant distributions for the deterministic logistic map, by looking at distribtuions invariant under $f_{\lambda}(x)=\lambda x(1-x)$.
+If, for all $x$, $Pr\{X\leq x\}=Pr\{f(X)\leq x\}$ then we say that the distribution of $X$ is invariant under $f$.
+In this way, we can discuss invariant distributions for the deterministic logistic map by looking at distributions invariant under $f_{\lambda}(x)=\lambda x(1-x)$.
 In [21], Ulam and von Neumann proved that the continuous invariant distribution for the deterministic logistic map with $\lambda=4$ is a beta distribution with $\alpha=\beta=0.5$.
-This function is symmetric on the interval (0,1), and goes to infinity at the endpoints of its support.
-Thus, the invariant distribution when $\lambda=4$ is given by the function
+The density for this distribution is symmetric on the interval (0,1), and goes to infinity at the endpoints of its support.
+To be specific, the invariant distribution when $\lambda=4$ is given by the density function
 $f^{*}(x)=\begin{cases}\frac{1}{\pi\sqrt{x(1-x)}}&0<x<1\\ 0&\hbox{ else}\end{cases}$ (6)
@@
 Figure 1: Histogram of the beta distribution with $\alpha=\beta=0.5$.
-The solid line shows the theoretical distribution given in Equation (6); the bars are a histogram generated by sampling 100,000 initial values uniformly on $(0,1)$
+The solid line shows the theoretical distribution given in Equation (6); the bars are a histogram generated by sampling 100,000 initial values uniformly on $(0,1)$ and then iterating 20 times.
 However, invariant densities for other values of $\lambda$ are not, in general, symmetric.
@@
 They also show that these invariant distributions were not symmetric.
-Figure 2 shows numerical approximations to theses deterministic invariant densities for $\lambda=3.87$ and $\lambda=3.95$ These numerical approximations are histograms of 100,000 sampled initial values iterated 2,000 times.
+Figure 2 shows numerical approximations to theses deterministic invariant densities for $\lambda=3.87$ and $\lambda=3.95$ These numerical approximations are histograms of 100,000 sampled initial values iterated 20 times.
 They indicate the singularities in the theoretical densities described in Remark 5 of [22].
@@
     A uniform distribution on (0,1) (Figure 3, center).
-In each case, after 2,000 iterations, we obtained a histogram very similar to that presented in Figure 4: a bimodal, asymmetric distribution that appears to be supported throughout the unit interval, without singularities (poles) in the interior.
-This is noteworthy, particularly because we simulated using both continuous and discrete starting distributions.
+In each case, after several iterations, we obtained a histogram very similar to that presented in Figure 4: a bimodal, asymmetric distribution that appears to be supported throughout the unit interval, without singularities (poles) in the interior.
+This is noteworthy, particularly because we simulated using both continuous and discrete initial distributions.
 We examined distributions that were left-skewed, right-skewed, and symmetric.
-Each simulated initial distribution appears to converge to the same distribution after 2,000 iterations, aligning with the global asymptotic stability of the invariant distribution established theoretically earlier in this paper.
+Each simulated initial distribution appears to converge to the same distribution, aligning with the global asymptotic stability of the invariant distribution established theoretically earlier in this paper.
 We also note that, while the distribution demonstrates peaks toward the sides of the interval, it does not exhibit the poles at the boundary of the interval in the same manner as the invariant distributions for the deterministic maps.
 However, this seems reasonable, as the interval of support of the invariant distribution is dependent on $\lambda$, as discussed above, and relatively fewer values of $\lambda$ will map points into ranges closer to 0 and 1, so it makes sense that the distribution will decrease near the boundary of $[0,1]$.
-It is also of note that even as the width of the interval that the distribution of $\lambda$ is supported on narrows, the invariant distribution maintains this property of decreasing near the boundary of support.
+It is also of note that even as the width of the interval on which $\lambda$ is supported narrows, the invariant distribution maintains this property of decreasing near the boundary of support.
 For instance, we ran simulations where $\lambda$ was sampled uniformly on the interval $(3.87,9)$ and again on the interval $(3.87,3.935).$ Both invariant distributions exhibited this decrease toward the edge of the boundary of support (which, as expected, is not the entirety of $(0,1)$.
@@
 Figure 6 shows the evolution of the simulated invariant density after various numbers of iterations, starting from a uniform distribution.
+In this set of simulations, the empirical distributions after 5, 20 and 200 iterations are quite close to each other, and the distribution after 20 iterations is almost indistinguishable from the distribution after 200 iterations.
+This justifies our use of only 20 iterates in the other simulations presented here.
 Refer to caption Refer to caption Refer to caption
 Figure 3: Starting distributions of 100,000 starting values sampled on the set $\{0.11,0.33,0.55,0.6,0.78\}$ with probabilities $\{0.196,0.140,0.233,0.322,0.107\}$ respectively (left), the uniform distribution on (0,1) (middle), and the truncated exponential distribution with parameter $r=1.25$ (right).
-Figure 4: Empirical distribution of $\{X_{n}\}$ after $n=2,000$ iterations, where 100,000 initial values were sampled from a truncated exponential distribution with parameter $r=1.25$ as in Figure 3.
+Figure 4: Empirical distribution of $\{X_{n}\}$ after $n=20$ iterations, where 100,000 initial values were sampled from a truncated exponential distribution with parameter $r=1.25$ as in Figure 3.
 Similar results were obtained using other initial distributions (results not shown).
 Figure 5: The simulated invariant distribution of $\{X_{n}\}$ for $\lambda\sim U(3.87,3.9)$ (left) and $\lambda\sim U(3.87,3.935)$ (right), shown after 20 iterations.
-Figure 6: The densities for a simulation of 100,000 values, initially sampled randomly from a uniform distribution on (0,1), after various numbers of iterations.
+Figure 6: The densities for a simulation of 100,000 values, initially sampled randomly from a uniform distribution on (0,1), after various numbers of iterations where $\lambda$ is uniformly randomly sampled from $(3.87,4)$ each iteration.
 5 Conclusion
-In conclusion, we have illustrated that the stochastic logistic map, characterized by parameter values uniformly distributed in the chaotic regime, exhibits Harris irreducibility.
+In conclusion, we have illustrated that the stochastic logistic map, characterized by parameter values uniformly distributed in the chaotic regime, is irreducible.
 Consequently, the previously established invariant distribution stands as the unique invariant distribution.
 This result also implies that all initial distributions converge to the invariant distribution in the Cesàro sense.
-Moreover, we have demonstrated the strong aperiodicity of the stochastic logistic map, indicating that initial distributions converge to the invariant distribution with respect to the total variation distance—a more robust and rigorous result.
+Moreover, we have demonstrated the strong aperiodicity of the stochastic logistic map, indicating that initial distributions converge to the invariant distribution with respect to the total variation distance—a more robust result.
 We conducted simulations of the stochastic logistic map to visualize the invariant distribution.
 Remarkably, the empirically derived invariant distribution exhibits qualitative distinctions when compared to the invariant distributions of the deterministic logistic map for values of $\lambda$ in the chaotic regime.
-In particular, the simulated distribution does not appear to have a pole at $x=0$, which contrasts with the work in [21], which demonstrates that the deterministic invariant density will have a pole at the boundary of its support.
+In particular, the simulated distribution does not appear to have a pole at $x=0$.
+This contrasts with the work in [22], which demonstrates that the deterministic invariant density will have a pole at the boundary of its support.
 It should be noted that this work does not seem to depend on the $\lambda$ values being uniformly distributed on (3.87,4) and indeed, it appears that the results will hold for any absolutely continuous distribution on (3.87,4).

### v4

@@
 This will be important in the proofs in Section 3.
-Let $\mathcal{B}(S)$ denote the Borel $\sigma$-algebra on $S$. We consider the logistic Markov process
+We consider the logistic Markov process
 $X_{n+1}=\lambda_{n+1}X_{n}(1-X_{n}),$ (1)
@@
 If $P\mu(A)=\mu(A)$ for all $A\in{\cal B}(S)$ then $\mu$ is an invariant measure under the operator $P$.
-We can iterate the operator $P$ a positive number $m$ times, which we denote by $P^{m}$: if $X_{0}$ has distribution $\mu$ then $X_{m}$ has distribution $P^{m}\mu$.
-This leads us to the analog of stability in the stochastic context.
-Definition 2.
-A sequence of probability measures $P^{n}(n\geq 1)$ on $S$ converges weakly to a probability measure $P$ on $S$ if, for any continuous, bounded function $h:S\mapsto S$,
-[…]
-In the deterministic case, we know that, if the iterates of $x_{0}$ under a continuous function, $f$, converge to a point, $x$, then $x$ must be a fixed point of $f$, and this result follows from the continuity of $f$.
-We can get a similar result in the stochastic setting, but we first need to define a notion of continuity.
-Suppose we have a Markov process on $S$ with transition probabilities given by $p(x,dy)$.
-Definition 3.
-The transition probability $p(x,dy)$ is weakly continuous or has the Feller property if, for every continuous and bounded function $h:S\mapsto S$, the function
-$x\mapsto\int_{S}h(y)p(x,dy)$
-is continuous.
-Using weak continuity we get the following result ([14], Theorem 11.1 of Chapter 2).
-Proposition 1.
-Suppose $p(x,dy)$ is a transition probability on $S$ with the Feller property.
-  1. 1.
-    If there is an $x\in S$, a sequence of integers $1\leq n_{1}<n_{2}<\dots$ and a probability measure $\pi_{x}$ on $S$ such that
-    $\frac{1}{n_{k}}\sum_{m=0}^{n_{k}-1}p^{m}(x,dy)\hbox{ converges weakly to }\pi_{x}\hbox{ as }k\rightarrow\infty$
-    then $\pi_{x}$ is an invariant measure for $p$.
-  2. 2.
-    If, for all $x\in S$, there is a sequence of integers $\{n_{i}\}$ such that the limit above exists and is independent of $x$ then $\pi=\pi_{x}$ is the unique invariant probability for $p$.
-We can state this proposition in terms of the operator $P$ given in Equation (2).
-Suppose a process has the Feller property and we have a measure $\mu$ on $S$.
-If we have weak convergence in terms of the average to a measure, $\pi$:
-$\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{m=0}^{n-1}P^{m}\mu(A){\rightarrow}\pi(A)$
-then $\pi$ is an invariant measure under the operator $P$, i.e.
-$P\pi(A)=\pi(A)$ for $A\in{\cal B}(S)$.
-Furthermore, if the condition in Item 2 of Proposition 1 holds, then this invariant measure is unique.
 What we would like is the asymptotic stability of the (unique) invariant measure, $\pi$. In other words, for any probability measure $\mu$ on $S$, when is it the case that
@@
   2. (ii)
-    The Markov process is $\phi$-recurrent if, for each $x\in S$ and $A\in\mathcal{B}(S)$ such that $\phi(A)>0$, we have
+    The Markov process is $\nu$-recurrent if, for each $x\in S$ and $A\in\mathcal{B}(S)$ such that $\nu(A)>0$, we have
     $P(\eta_{A}<\infty|X_{0}=x)=1,\quad\hbox{ where }\;\eta_{A}=\inf\{n\geq 1:X_{n}\in A\}$
   3. (iii)
-    If the Markov process is both $\nu$-irreducible and $\phi$-recurrent, then it is Harris recurrent.
+    If the Markov process is both $\nu$-irreducible and $\nu$-recurrent, then it is Harris ($\nu$-) recurrent.
   4. (iv)
@@
     $\sup\{E(\eta_{A_{0}}|X_{0}=x):x\in A_{0}\}<\infty$
-    for some $\phi$-recurrent set $A_{0}$, then the process is said to be positive Harris recurrent.
-Note: the set $A_{0}$ is $\phi$-recurrent if $\phi(A_{0})>0$ and it is recurrent, i.e. $P(\eta_{A_{0}}<\infty|X_{0}=x)=1$ $\forall x\in{\mathcal{B}}(S)$.
+    for some $\nu$-recurrent set $A_{0}$, then the process is said to be positive Harris ($\nu$) recurrent.
 In other words: a Markov process is $\nu$-irreducible if there is a positive probability that every element in the state space eventually gets mapped to any set of positive $\nu$ measure.
-It is $\phi$-recurrent if, with probability 1, it visits every set of positive $\phi$-measure in a finite number of steps.
-The Krylov-Bogolubov Theorem [3] tells us that a stochastic dynamical system obtained from a continuous function on a compact subset of ${\mathbb{R}}^{d}$ with independent parametric noise always has an invariant distribution:
-Theorem 1 (Krylov-Bogolubov).
-A Feller continuous, stationary Markov process on a compact subset of ${\mathbb{R}}^{d}$ has an invariant distribution.
-The theorem applies to more general settings, but this more specific version is all we need here.
-Irreducibility can then be used to show that this invariant distribution is unique using the following theorem:
+It is $\nu$-recurrent if, with probability 1, it visits every set of positive $\nu$-measure in a finite number of steps.
+Definition 3.
+The set $A_{0}$ is $\nu$-recurrent if $\nu(A_{0})>0$ and it is recurrent, i.e. $P(\eta_{A_{0}}<\infty|X_{0}=x)=1$ $\forall x\in{\mathcal{B}}(S)$.
+Irreducibility can then be used to show that the invariant distribution is unique using the following theorem:
 Theorem 1.
-[[14], Theorem C9.3 in Chapter 2] If a Markov process is irreducible with respect to some $\sigma$-finite measure $\nu$ on $S$ and has an invariant probability $\mu^{*}$, then the process is positive Harris recurrent, $\mu^{*}$ is the unique invariant probability and, for any probability measure $\mu$ on S:
+[[14], Theorem C9.3 in Chapter 2] If a Markov process is $\nu$-irreducible with respect to some $\sigma$-finite measure $\nu$ on $S$ and has an invariant probability $\mu^{*}$, then the process is positive Harris ($\mu^{*}$-) recurrent, $\mu^{*}$ is the unique invariant probability and, for any probability measure $\mu$ on S:
 $\lim_{n\rightarrow\infty}\sup_{A\in\mathcal{B}(S)}\left|\frac{1}{n}\sum_{m=1}^{n}P^{m}\mu(A)-\mu^{*}(A)\right|=0$
@@
 the average of the sequence of measures converges to $\mu^{*}$.
-We also get positive Harris recurrence from this result.
+Note that a result from [15] (Proposition 10.1.2) states that the invariant measure $\mu^{*}$ dominates the irreducible measure $\nu$ (that is, every set of positive $\nu$ measure also has positive $\mu^{*}$ measure), so Theorem 1 also implies positive Harris recurrence with respect to $\nu$.
 To get asymptotic stability of the invariant measure without this average, we need one more condition: aperiodicity.
 Definition 4.
-A Markov process is said to be strongly aperiodic with respect to a $\sigma$-finite probability measure $\phi$ on $S$ if there is a set $A_{0}\subset S$ with $\phi(A_{0})=1$ and an $\alpha>0$ such that one has
+A Markov process is said to be strongly aperiodic with respect to a $\sigma$-finite probability measure $\nu$ on $S$ if there is a set $A_{0}\subset S$ with $\nu(A_{0})=1$ and an $c>0$ such that one has
   1. 1.
-    for all $x\in S$, $P(\eta_{A_{0}}<\infty|X_{0}=x)=1$, i.e. the process is $A_{0}$-recurrent;
+    for all $x\in S$, $P(\eta_{A_{0}}<\infty|X_{0}=x)=1$, i.e. $A_{0}$ is recurrent;
   2. 2.
-    for all $x\in A_{0}$, $A$ a measurable subset of $S$, $p(x,A)\geq\alpha\phi(A)$.
+    for all $x\in A_{0}$, $A$ a measurable subset of $S$, $p(x,A)\geq c\nu(A)$.
 Let $\{X_{n}\}$ be a Markov process on a state space $S$, and $\mathcal{B}$} the Borel subsets of $S$.
 Theorem 2 ([16], Corollary 6.7).
-Suppose $\{X_{n}\}$ is a positive Harris recurrent and aperiodic Markov process with stationary distribution $\pi$. Then for any initial distribution, $\mu$, of $X_{0}$:
+Suppose $\{X_{n}\}$ is a positive Harris recurrent and strongly aperiodic Markov process with respect to some nonzero $\sigma$-finite measure $\nu$, with stationary distribution $\pi$.
+Then for any initial distribution, $\mu$, of $X_{0}$:
 $\lim_{n\rightarrow\infty}\delta(P^{n}\mu,\pi)=0$
@@
 $\delta(P,Q)=\sup_{B\in\mathcal{B}}|P(B)-Q(B)|.$
-We note that the measures used in the definitions of irreducibility and aperiodicity could be different.
-(This fact will be used later in the proof of Theorem 5.) To see that this is the case, consider the following.
-Aperiodicity with respect to the measure $\nu$ implies that there is an $\alpha>0$ such that
-$p(x,B)\geq\alpha\nu(B)\,\Rightarrow\,p(x,B)>0\;\hbox{ if }\nu(B)>0$
-Thus, the process is $\nu$-irreducible. Therefore, in the statement of Theorem 3, we do not need to specify the measures used to show either irreducibility or aperiodicity.
 As an example of the ideas presented in this section, let us consider a version of the stochastic logistic map as examined in [17].
@@
 If $1<\alpha<\beta\leq 2$ then there is a unique invariant probability on (0,1), and this probability is nonatomic.
-If $2<\alpha<\beta<1+\sqrt{5}$ and $\alpha\in[8/\beta(4-\beta),\beta)$, then there exists a unique invariant probability on (0,1) which is nonatomic and has its support contained in $[1/2,(1+\sqrt{5})/4]$.
+If $2<\alpha<\beta<1+\sqrt{5}$ and $\alpha\in[8/\beta(4-\beta),\beta)$, then there exists a unique invariant probability on (0,1) which is nonatomic and has its support contained in $1/2,(1+\sqrt{5})/4]$.
 3 Proof of the asymptotic stability of the invariant measure in the chaotic regime.
@@
 In this section we use these results to prove that, when parameters are supported in the chaotic regime, the random logistic map has a unique asymptotically stable invariant measure.
-Recall that the deterministic logistic family: $f(x,\lambda)=f_{\lambda}(x)=\lambda x(1-x)$ undergoes a period-doubling cascade as $\lambda$ increases from 1 to $\lambda_{2\infty}\approx$ 3.56995.
+Recall that the deterministic logistic family: $f(x,\lambda)=f_{\lambda}(x)=\lambda x(1-x)$ undergoes a period-doubling cascade as $\lambda$ increases from 1 to $\lambda_{2^{\infty}}\approx$ 3.56995.
 Note that here iteration occurs only with respect to $x$; that is $f^{(m)}(x,\lambda)=f(f^{m-1}(x,\lambda),\lambda)$, where $\lambda$ remains fixed.
@@
 Interspersed between the values of $\lambda$ with stable periodic orbits are a set of large measure of parameter values for which there are no stable periodic orbits, and almost all initial conditions give rise to orbits with no finite period.
-Thus, $\lambda_{2\infty}$ marks the “onset of chaos”.
-As $\lambda$ increases beyond $1+\sqrt{8}$, the family of maps again experiences a period-doubling cascade, with stable orbits of period 6, 12, 24, 48, …, culminating at $\lambda_{3\infty}\approx 3.8495$.
+Thus, $\lambda_{2^{\infty}}$ marks the “onset of chaos”.
+As $\lambda$ increases beyond $1+\sqrt{8}$, the family of maps again experiences a period-doubling cascade, with stable orbits of period 6, 12, 24, 48, …, culminating at $\lambda_{3^{\infty}}\approx 3.8495$.
 When $\lambda$ is greater than this value, the maps exhibit chaotic behavior on a Cantor set of parameter values, with three “chaotic bands” of orbits merging into one chaotic band at $\lambda\approx 3.857$.
@@
 To make sure we are well within the parameter range for which there is generically a single band chaotic attractor, we restrict $\lambda\in(3.87,4)$.
-The logistic family is well-studied in the literature, and more details on its dynamics can be found in, for example, [17]
-In [18], Athreya and Dai consider the stochastic logistic map and demonstrate that the Markov process $\{X_{n}\}$ is Feller continuous, which guarantees the existence of an invariant distribution (but not its stability) by Theorem 1.
-They also give conditions for convergence of the distribution of the $X_{i}$ in the sense of Césaro which differ somewhat from the hypotheses of Theorem 2.
-Here we choose to follow Bhattacharya and Majumbdar ([14]) and Nummelin ([15]) by showing the irreducibility and aperiodicity of the Markov process in the chaotic regime.
-This, along with the existence of an invariant measure, gives the asymptotic stability of the invariant measure by Theorem 3.
-Lemma 1 and its generalization Lemma 2 build toward Theorem 4, which shows that the Markov process is irreducible.
-Theorem 5 shows that the process is strongly aperiodic and thus, since it has an invariant measure by Theorem 1, the process has a unique asymptotically stable invariant measure.
+The logistic family is well-studied in the literature, and more details on its dynamics can be found in, for example, [18].
+We consider the stochastic logistic map, i.e. the Markov process given by
+$X_{n+1}=\lambda_{n+1}X_{n}(1-X_{n})$ (3)
+where $\{\lambda\}_{i\geq 1}$ are i.i.d.
+random variables distributed uniformly on (3.87,4).
+The initial state $X_{0}$ is a random variable taking values in (0,1), with distribution independent of the $\lambda$’s.
+In ([19] Theorem 2), Athreya and Dai demonstrate that the Markov process $\{X_{n}\}$ given by Equation (3) has an invariant distribution.
+We structure this section as follows: we first show the irreducibility of the Markov process in the chaotic regime using Lemmas 1 and 2, which together with the existence of the invariant measure gives positive Harris recurrence, by Theorem 1.
+This, along with strong aperiodicity, gives the asymptotic stability of the invariant measure by Theorem 2.
+The following lemma uses the existence of a dense orbit and continuity of the iterated functions to show that every $x\in(0,1)$ has a positive probability of entering any open set in finite time.
+This lemma will be used to prove the irreducibility of $\{X_{n}\}$.
 Lemma 1.
-Let $\lambda\sim\mathcal{U}(3.87,4)$, $\alpha$ a fixed element in $(3.87,4)$, and $x^{*}\in(0,1)$ such that $x^{*}=\alpha x^{*}(1-x^{*})$.
-(Equivalently, $x^{*}=1-\frac{1}{\alpha}$.) Then there exists $\eta>0$ such that for all $x\in I=(x^{*}-\eta,x^{*}+\eta)$ and $A\subset I$ with $m(A)>0$, then
-$p(x,A)=\hbox{Prob}\{X_{1}\in A\,|\,X_{0}=x\}>0.$
-where $m$ is Lebesgue measure.
-That is, there is a neighborhood $I$ around the fixed point for which every $x\in I$ has a positive probability of visiting each subset of $I$ of positive Lebesgue measure.
+Let $A$ be a nonempty open subset of $(0,1).$ For all $x\in(0,1)$ there is an $M$ such that $p^{M}(x,A)=Prob(X_{M}\in A|X_{0}=x)>0$.
 Proof.
-Fix $\alpha\in(3.87,4)$, and let $x^{*}\in(0,1)$ be such that $\alpha x^{*}(1-x^{*})=x^{*}$.
-Notice that for any $\eta>0$ and $A\subset I\equiv(x^{*}-\eta,x^{*}+\eta)$ with $m(A)>0$
-[…]
-and
-[…]
-Since $h(x)=x(1-x)$ is continuous at $x^{*}$ and $\alpha h(x^{*})=x^{*}$, it follows that there exists $\eta^{*}>0$ such that $a(\eta^{*})<4$ and $b(\eta^{*})>3.87$.
-Let $A_{\eta^{*}}\subset I_{\eta^{*}}=(x^{*}-\eta^{*},x^{*}+\eta^{*})$ be a set of positive Lebesgue measure associated with this particular value of $\eta$.
-Then, for any $x\in S$, ${\frac{A_{\eta^{*}}}{x(1-x)}}\subset(3.87,4)$, and since $A_{\eta^{*}}$ has positive measure, so does $\frac{A_{\eta^{*}}}{x(1-x)}$, and therefore $p(x,A_{\eta^{*}})>0$.
-Remark 1.
-Note that while we take the distribution of $\lambda$ on $(3.87,4)$ to be uniform, the same result will hold if we replace the uniform distribution with a distribution that has non-vanishing absolutely continuous part on this interval.
-The next lemma generalizes Lemma 1 to the case in which the map has a stable periodic orbit of any period.
+It is known [20] that there is a value $\alpha\in(3.87,4)$ such that $f(x,\alpha)=\alpha x(1-x)$ exhibits topological transitivity; that is $f(x,\alpha)$ contains an orbit that is dense in $0,\alpha/4)$.
+Call this dense orbit $\Gamma.$ Let $x\in(0,1)$ Note then that, for all $x\in(0,1)$, the set $\{\lambda x(1-x):3.87<\lambda<4\}$ is an open interval intersecting $(0,\alpha/4)$, and therefore must contain a point on $\Gamma.$ Therefore, there exists $\beta\in(3.87,4)$ such that $y=f(x,\beta)$ is on $\Gamma$.
+Note then, that since the orbit of $y$ is dense for parameter $\alpha$, there exists $M$ such that $f^{M}(y,\alpha)\in A$.
+Then, $f^{M}(f(x,\beta),\alpha)\in A$.
+Note further that $f(\ldots,f(x,\lambda_{1}),\cdots,\lambda_{M+1})$ is a polynomial in $M+2$ variables and is thus continuous everywhere.
+Therefore, since $f^{M}(f(x,\beta),\alpha)\in A$, and $A$ is an open set, there is a set of positive normalized Lebesgue measure $U$ in $(3.87,4)^{M+1}$ containing $(\beta,\underbrace{\alpha,\;\cdots\;,\alpha}_{\hbox{$M$ times}})$ such that for all $(\lambda_{1},\cdots,\lambda_{M+1})\in U$, $f(\ldots,f(x,\lambda_{1}),\cdots,\lambda_{M+1})\in A$.
+This means that
+$Prob(X_{M+1}\in A|X_{0}=x)>0.$
+The following lemma is used later to show strong aperiodicity as well as irreducibility.
 Lemma 2.
-As before, let $\lambda_{i}\sim\mathcal{U}(3.87,4)$ for all $i\geq 1$.
-Suppose $\alpha$ is a parameter value in $(3.87,4)$ with an $M$-periodic point, $x^{*}$, so that $f^{(M)}(x^{*},\alpha)=x^{*}$.
-Then there is an $\eta>0$ such that, for any $x\in I=(x^{*}-\eta,x^{*}+\eta)$ and $A\subset I$ with $m(A)>0$, we have
-$p^{M}(x,A)=Prob(X_{M}\in A|X_{0}=x)>0$
-Proof.
-Since $43.87<\alpha<$ 4 we can choose an $\eta_{1}$ small enough so that
-$43.87<\left({\frac{x^{*}-\eta_{1}}{x^{*}+\alpha\eta_{1}}}\right)\alpha<\left({\frac{x^{*}+\eta_{1}}{x^{*}-\alpha\eta_{1}}}\right)\alpha<$ 4
-because
-$\lim_{\eta_{1}\rightarrow 0}\frac{x^{*}-\eta_{1}}{x^{*}+\alpha\eta_{1}}=\lim_{\eta_{1}\rightarrow 0}\frac{x^{*}+\eta_{1}}{x^{*}-\alpha\eta_{1}}=1$
-Denote points in the periodic orbit of $x^{*}$ as $x^{*}_{j}=f^{(j)}(x^{*},\alpha)$.
-Since $x^{*}$ has period $M$, $f(x^{*}_{M-1},\alpha)=\alpha x^{*}_{M-1}(1-x^{*}_{M-1})=x^{*}$.
-Since $f(\cdot,\cdot)$ is continuous in both variables:
-  1. 1.
-    there is an $\eta_{2}>0$ such that, if $|y-x^{*}_{M-1}|<\eta_{2}$ then
-    $|f(y,\alpha)-x^{*}|<\eta_{1}\alpha\Leftrightarrow\left|y(1-y)-\frac{x^{*}}{\alpha}\right|<\eta_{1}\,;$
-  2. 2.
-    given $\eta_{2}$ as in the previous statement, $\exists\,\eta_{3}$ so that $|x-x^{*}|<\eta_{3}$ and $|\lambda_{i}-\alpha|<\eta_{3}$ for $1\leq i<M\Rightarrow|X_{M-1}-x^{*}_{M-1}|<\eta_{2}$ where $X_{j+1}=\lambda_{j+1}X_{j}(1-X_{j})$ and $X_{0}=x$.
-In other words, if the random logistic map starts close to $x^{*}$, and if all of the parameter values are close to $\alpha$, then the random orbit returns close to $x^{*}$ after $M$ iterates.
-Let $0<\eta<\min\{\eta_{1},\eta_{3}\}$. Then, if $|\lambda_{i}-\alpha|<\eta$ for $1\leq i<M$, and $A\subset I=(p-\eta,p+\eta)$, with $X_{0}\in I$, we have:
-$\frac{A}{X_{M-1}(1-X_{M-1})}\subset\left({\frac{p-\eta_{1}}{\frac{p}{\alpha}+\eta_{1}},\frac{p+\eta_{1}}{\frac{p}{\alpha}-\eta_{1}}}\right)\subset(3.87,4)$
-by our original assumption on $\eta_{1}$.
-Since $A$ has positive measure, so does ${\frac{A}{X_{M-1}(1-X_{M-1})}}$.
-To help clarify the notation, let $C$ be the event that $\lambda_{i}\in(\alpha-\eta,\alpha+\eta)$ for $1\leq i<M$.
-What we have shown is that
-$Prob(X_{M}\in A|X_{M-1},C)=Prob\left({\lambda_{M}\in\frac{A}{X_{M-1}(1-X_{M-1})}|X_{M-1},C}\right)>0.$ (3)
-Assume that $X_{0}=x\in I$, and recall that $\lambda_{M}$ is independent of $X_{0},X_{1},\dots X_{M-1}$. Then by the law of total probability:
-$Prob(X_{M}\in A)\geq Prob((X_{M}\in A)\cap C)=E_{x}[Prob((X_{M}\in A)\cap C)|(X_{M-1}=x)]$ (4)
-where $E_{x}$ is the expected value with respect to the distribution of $X_{M-1}$. Focusing on what’s inside the expected value:
-$Prob\left({(X_{M}\in A)\cap C|X_{M-1}}\right)=Prob(X_{M}\in A|X_{M-1},C)Prob(C)$
-Since the $\lambda$’s are independent of $X_{i}$ and $Prob(|\lambda_{i}-\alpha|<\eta)>0$, we have by Equation (3)
-$Prob\left({(X_{M}\in A)\cap C|X_{M-1}}\right)=Prob(X_{M}\in A|X_{M-1},C)\left[{Prob(|\lambda_{i}-\alpha|<\eta)}\right]^{M-1}>0$
-The expected value of a strictly positive random variable is positive, so we conclude, using Equation (4) that, if $X_{0}\in I$ then $Prob(X_{M}\in A)>0$, as claimed. ∎
-Theorem 4.
-As above, let $S=[0,1]$.
+Let $S=(0,1)$.
 $X_{n+1}=\lambda_{n+1}X_{n}(1-X_{n}),\,\,\,n\geq 0$
 with $\{\lambda_{n}\}_{n\geq 1}$ being i.i.d., $\lambda_{i}\sim\mathcal{U}(3.87,4)$, and $X_{0}$ an independent $S$-valued r.v.
-Then $\{X_{n}\}$ is irreducible with respect to the Lebesgue measure $m$ on $S$ conditioned on $(x^{*}-\eta,x^{*}+\eta)$.
-Proof.
-By a result of Graczyk and Swiatek [19], the set of all $\lambda$ in $(0,4)$ such that $f(\cdot,\lambda)$ has a stable periodic orbit is dense.
-Thus, there is an $\alpha\in(3.87,4)$ and $10<x^{*}<$ 1 such that $x^{*}$ is a stable periodic point for $f(x,\alpha)=\alpha x(1-x)$.
-Let $K=\{x\in(0,1)\,\,|\,\,\omega_{\alpha}(x)=\gamma(x^{*})\}$, where $\omega_{\alpha}$ is the limit set of the orbit of $x$ under $f(\cdot,\alpha)$ and $\gamma(x^{*})$ is the orbit of the stable periodic point $x^{*}$.
-In [20] (Chapter V, Theorem 1.3) de Melo and van Strien have shown that $m(K)=1$ for unimodal maps, where $m$ is Lebesgue measure.
-Thus let $I=(x^{*}-\eta,x^{*}+\eta)$ as given by Lemma 2.
-Then, for all $x\in K$, $f^{(n)}(x,\alpha)\in I$ for some large $n$.
-Thus, by continuity of $f$,
-$Prob(X_{n}\in I\text{ for some }n\geq 1|X_{0}=x)>0$ (5)
-for all $x\in K$.
-We will next show that for any $x\in(0,1)$, there is a positive probability that $x$ is mapped to $K$ in one iterate.
-In fact, we will show that $p(x,K)=Prob(X_{1}\in K|X_{0}=x)=1$.
-Fix $x\in(0,1)$ and let $K^{c}$ be the complement of $K$ in $[0,1]$.
-Since $\lambda$ is uniformly distributed on $(3.87,4)$ and $m(K^{c})=0$ we have
-$p(x,K)$ $=$ $Prob(\lambda\,|\,\lambda x(1-x)\in K)$
-$=$ $Prob\left({\lambda\in\frac{K}{x(1-x)}}\right)$
-$=$ $Prob\left({(3.87,4)\cap\frac{K}{x(1-x)}}\right)$
-$=$ $1-Prob\left({(3.87,4)\cap\frac{K^{c}}{x(1-x)}}\right)$
-$\geq$ $1\displaystyle 1-Prob\left({\frac{K^{c}}{x(1-x)}}\right)=$ 1
-Therefore, $p(x,K)>0$, establishing irreducibility. ∎
-Using Theorems 1 and 2, we now have that, since the process is Feller continuous by [18], the process $\{X_{i}\}$ has a unique invariant measure that is stable in the sense of Césaro.
-We conclude this section by demonstrating that the process is strongly aperiodic, and thus the invariant distribution is asymptotically stable with respect to the total variation norm.
-Theorem 5.
-Let $S=[0,1]$.
-$X_{n+1}=\lambda_{n+1}X_{n}(1-X_{n}),\,\,\,n\geq 0$
-with $\{\lambda_{n}\}_{n\geq 1}$ being i.i.d., $\lambda_{i}\sim\mathcal{U}(3.87,4)$, and $X_{0}$ an independent $S$-valued r.v.
-Then $\{X_{n}\}$ is strongly aperiodic, and thus has an asymptotically stable invariant distribution.
+Let $\lambda_{min}=3.87$, and let $A_{0}=(1-\frac{1}{\lambda_{min}},1-\frac{1}{4})$ (that is, $A_{0}$ is the set of fixed points for the deterministic logistic maps corresponding to the possible parameter $\lambda$ values).
+Let $\phi$ be normalized Lebesgue measure on $A_{0}$.
+Then there exists $c>0$ such that for all $x\in A_{0}$, and $A$ a measurable subset of $(0,1)$, we have $p(x,A)\geq c\phi(A).$
 Proof.
@@
 Let $L_{1}$ be the length of $A_{0}$ (that is, $L_{1}=\frac{1}{\lambda_{min}}-\frac{1}{4}$ ) and let $L_{2}$ be the length of interval $(\lambda_{min},4)$ (that is, $L_{2}=4-\lambda_{\min}$).
-Let $\nu$ be normalized Lebesgue measure on $A_{0}$.
+Let $\phi$ be normalized Lebesgue measure on $A_{0}$.
 Then $A_{0}$, by definition, has measure 1, and for any measurable set $A$ :
-[…]
+$\phi(A)=\frac{1}{L_{1}}\int_{y\in(A\cap A_{0})}dy$
 Fix $x\in A_{0}$ and $A\subset S$.
-We want to show that $p(x,A)\geq c\,\nu(A)$ for some positive constant, $c$ (Definition 5).
-Since $\nu(A)=\nu(A\cap A_{0})$ it suffices to demonstrate the above result for sets $A\subset A_{0}$.
-Notice for any $A\subset A_{0}$ and with $\nu(A)>0$, and $x\in A_{0}$,
-[…]
+We want to show that $p(x,A)\geq c\,\phi(A)$ for some positive constant, $c$ (Definition 4).
+Since $\phi(A)=\phi(A\cap A_{0})$ it suffices to demonstrate the above result for sets $A\subset A_{0}$.
+Notice for any $A\subset A_{0}$ and with $\phi(A)>0$, and $x\in A_{0}$,
+$\underset{x}{\sup}\left\{\frac{A}{x(1-x)}\right\}<\frac{1-\frac{1}{4}}{\frac{1}{4}\left(1-\frac{1}{4}\right)}=4$
 and
-[…]
+$\underset{x}{\inf}\left\{\frac{A}{x(1-x)}\right\}>\frac{1-\frac{1}{\lambda_{\min}}}{\frac{1}{\lambda_{\min}}\left(1-\frac{1}{\lambda_{\min}}\right)}=\lambda_{\min}=3.87$
 Therefore, for any $A\subset A_{0}$ and $x\in A_{0}$, $\frac{A}{x(1-x)}\subset(3.87,4).$
@@
 $p(x,A)$ $=$ $Prob\left(\lambda\in\frac{A}{x(1-x)}\right)$
-[…]
-[…]
-[…]
-$=$ $\frac{1}{L_{2}}\cdot\frac{1}{x(1-x)}\cdot L_{1}\nu(A)$
-$\geq$ $4\cdot\frac{L_{1}}{L_{2}}\nu(A),$
-since $x(1-x)\leq\frac{1}{4}$ for all $x\in[0,1]$.
-Thus, letting $\alpha=4\cdot\frac{L_{1}}{L_{2}}$ gives us the desired result.
-Therefore, the process $\{X_{n}\}$ is strongly aperiodic, and has an asymptotically stable invariant distribution.
-∎
+$=$ $\displaystyle\frac{1}{L_{2}}\int_{\frac{A}{x(1-x)}}d\lambda$
+$=$ $\frac{1}{L_{2}}\int_{A}\frac{1}{x(1-x)}dy$
+$=$ $\frac{1}{L_{2}}\frac{1}{x(1-x)}\int_{A}dy$
+$=$ $\frac{1}{L_{2}}\cdot\frac{1}{x(1-x)}\cdot L_{1}\phi(A)$
+$\geq$ $4\cdot\frac{L_{1}}{L_{2}}\phi(A)$
+$=$ $\frac{1}{\lambda_{\min}}\phi(A)\,$,
+since $x(1-x)\leq\frac{1}{4}$ for all $x\in[0,1]$. Thus, letting $c=\displaystyle{\frac{1}{\lambda_{\min}}}$ gives us the desired result. ∎
+Theorem 3.
+Let $S=(0,1)$, and set $A_{0}$ be defined as above. As before, define the Markov process on $S$:
+$X_{n+1}=\lambda_{n+1}X_{n}(1-X_{n}),\,\,\,n\geq 0$
+with $\{\lambda_{n}\}_{n\geq 1}$ being i.i.d., $\lambda_{i}\sim\mathcal{U}(3.87,4)$, and $X_{0}$ an independent $S$-valued r.v.
+Then $\{X_{n}\}_{n=0}^{\infty}$ is irreducible with respect to normalized Lebesgue measure on $A_{0}$.
+Proof.
+Let $x\in(0,1)$, and let $B$ be a subset of positive measure. We’d like to show that there is some $n\geq 1$ such that $p^{n}(x,B)=P(X_{n}\in B|X_{0}=x)>0.$ Note that
+$P(X_{n}\in B|X_{0}=x)=P(X_{n}\in B|X_{n-1}\in A_{0})*P(X_{n-1}\in A_{0}|X_{0}=x).$
+By Lemma 2, $P(X_{n}\in B|X_{n-1}\in A_{0})\geq c\,\phi(B)>0$, where $c=\frac{1}{\lambda_{min}}$.
+$P(X_{n-1}\in A_{0}|X_{0}=x)>0$ by Lemma 1, since $A_{0}$ is an open interval.
+Thus, $p^{n}(x,B)=P(X_{n}\in B|X_{0}=x)>0$ and $\{X_{n}\}_{n=0}^{\infty}$ is irreducible with respect to normalized Lebesgue measure on $A_{0}$, ∎
+By the existence of the invariant measure by [19] and Theorem 1, the process is positive Harris recurrent with respect to $\phi$, and the process $\{X_{i}\}$ has a unique invariant measure that is stable in the sense of Césaro.
+We conclude this section by demonstrating that the process is strongly aperiodic, and thus the invariant distribution is asymptotically stable with respect to the total variation norm.
+Theorem 4.
+Let $S=(0,1)$.
+$X_{n+1}=\lambda_{n+1}X_{n}(1-X_{n}),\,\,\,n\geq 0$
+with $\{\lambda_{n}\}_{n\geq 1}$ being i.i.d., $\lambda_{i}\sim\mathcal{U}(3.87,4)$, and $X_{0}$ an independent $S$-valued r.v.
+Then $\{X_{n}\}_{n=0}^{\infty}$ has an invariant distribution that is asymptotically stable with respect to the total variation norm.
+Proof.
+Because $\{X_{n}\}_{n=0}^{\infty}$ is positive Harris recurrent with respect to $\phi$, it is $\phi$-recurrent.
+This together with Lemma 2 gives that the process is strongly aperiodic with respect to $\phi$.
+By Theorem 2, this implies that $\{X_{n}\}$ has an invariant distribution that is asymptotically stable with respect to the total variation norm.
 4 Simulations and Discussion
@@
 In addition, Hall and Wolff observed that the invariant density will be supported on the interval $(\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}),\frac{1}{4}\lambda)$.
-This is due to the fact that any point in the interval $(\frac{1}{4}\lambda,1)$ will be mapped into the interval $(0,\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}))$, and any point in $(0,\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4})$ will be mapped to a point in $[\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}),\frac{1}{4}\lambda]$.
-Therefore, after finitely many iterations, any point $x\in(0,1)$ must be mapped into $[\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}),\frac{1}{4}\lambda]$, and so the invariant density must be supported on this subinterval.
+This is due to the fact that any point in the interval $(\frac{1}{4}\lambda,1)$ will be mapped into the interval $(0,\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}))$, and any point in $(0,\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4})$ will be mapped to a point in $\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}),\frac{1}{4}\lambda]$.
+Therefore, after finitely many iterations, any point $x\in(0,1)$ must be mapped into $\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}),\frac{1}{4}\lambda]$, and so the invariant density must be supported on this subinterval.
 Note that as $\lambda\rightarrow 4$, $\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4})\rightarrow 0$ and $\frac{1}{4}\lambda\rightarrow 1$, so the interval of support widens as $\lambda$ approaches 4.